### v2

# Reconstructing Subnational Labor Indicators in Colombia: An Integrated Machine and Deep Learning Approach

Abstract
This study proposes a unified multi-stage framework to reconstruct consistent monthly and annual labor indicators for all 33 Colombian departments from 1993 to 2025. The approach integrates temporal disaggregation, time-series splicing and interpolation, statistical learning, and institutional covariates to estimate seven key variables: employment, unemployment, labor force participation (PEA), inactivity, working-age population (PET), total population, and informality rate, including in regions without direct survey coverage. The framework enforces labor accounting identities, scales results to demographic projections, and aligns all estimates with national benchmarks to ensure internal coherence. Validation against official departmental GEIH aggregates and city-level informality data for the 23 metropolitan areas yields in-sample Mean Absolute Percentage Errors (MAPEs) below 2.3% across indicators, confirming strong predictive performance. To our knowledge, this is the first dataset to provide spatially exhaustive and temporally consistent monthly labor measures for Colombia. By incorporating both quantitative and qualitative dimensions of employment, the panel enhances the empirical foundation for analyzing long-term labor market dynamics, identifying regional disparities, and designing targeted policy interventions.
Keywords: Labor statistics; temporal disaggregation; machine learning; deep learning; Colombia
1 Introduction
Accurate, timely, and spatially disaggregated labor statistics are crucial for diagnosing employment dynamics, informing policy decisions, and advancing regional equity within decentralized governance frameworks OIT & CEPAL, (2022); OECD, (2021). However, in Colombia, structural constraints have resulted in significant gaps in subnational labor monitoring. As of 2025, the national household survey (GEIH, Gran Encuesta Integrada de Hogares) provides consistent indicators for only 23 of the 33 departments. Since December 2021, city-level series for 32 major urban areas have been released; however, these data do not directly correlate with departmental labor conditions in Amazonas, Arauca, Caquetá, Chocó, Guainía, Guaviare, Putumayo, San Andrés, Vaupés, or Vichada  [^1] Despite the fact that Gross Domestic Product (GDP) estimates are produced for all 33 departments,these capture production rather than labor outcomes.. This gap underscores the necessity for methodological efforts to align city and departmental labor statistics, which is the central focus of this study.
The limited availability of historical, department-level labor data constrains the study of long-term trends, cyclical fluctuations, and regional disparities. These limitations are particularly significant for analyzing informality, which constitutes a substantial portion of the Colombian workforce DANE, (2025). To address this gap, we introduce a unified estimation framework that reconstructs consistent monthly labor indicators for all 33 Colombian departments over the period 1993-2025. The framework encompasses standard labor aggregates and distinguishes between formal and informal employment  [^2] Subemployment and other highly disaggregated labor indicators are omitted due to insufficient and inconsistent information across time and departments. It integrates temporal disaggregation, demographic anchoring, statistical learning techniques, and the enforcement of labor accounting identities, with all estimates calibrated to national benchmarks.
The outcome is a demographically consistent and statistically coherent panel that aligns fully with official aggregates. For the first time, the resulting dataset provides comprehensive monthly coverage of seven core variables: employment, unemployment, inactivity, labor force (PEA, Población Económicamente Activa), working-age population (PET, Población en Edad de Trabajar), total population, and the informality rate. This enables systematic comparisons across various time periods and regions, including those prior to 2001, when no official high-frequency statistics were available.
[…]
Beyond its technical contribution, the reconstructed panel provides a robust empirical foundation to trace long-run regional labor trajectories, assess the heterogeneous effects of macroeconomic crises, and analyze structural asymmetries across departments. A composite Employment Quality Index (EQI) is incorporated to characterize territorial labor markets, explicitly penalizing configurations where high employment coexists with high informality. This multidimensional approach facilitates a more nuanced evaluation of convergence, divergence, and resilience within Colombia’s labor landscape.
To our knowledge, no existing dataset provides consistent monthly labor indicators, including informality, for all Colombian departments over such an extended period. Current official statistics remain fragmented, offering only partial subnational coverage. By addressing this gap, our dataset facilitates, for the first time, a systematic analysis of both the quantity and quality of employment across the national territory over more than three decades. This supports historical inquiry and evidence-based policy design.
The broader significance of this work extends to other countries facing similar statistical limitations. It illustrates how statistical learning methods can complement household surveys to expand the scope of labor monitoring in data-constrained contexts. Simultaneously, the results highlight a fundamental limitation: while model-based reconstruction can generate coherent and policy-relevant indicators, it cannot replace primary data collection in structurally underrepresented regions.
The remainder of this paper is structured as follows. Section 2 reviews the relevant literature on labor estimation, temporal disaggregation, and supervised prediction. Section 3 details the reconstruction pipeline. Section 4 presents the main empirical findings and validation metrics. Section 5 concludes and outlines future research directions.
2 Literature Review
This section reviews recent academic and technical contributions related to the estimation of labor market indicators under conditions of incomplete data coverage. The review is organized into five thematic areas: (i) applications of machine learning to labor statistics, (ii) temporal disaggregation of labor indicators, (iii) estimation under spatial gaps, (iv) regional segmentation and convergence, and (v) accounting consistency in predictive modeling. We prioritize studies conducted in Colombia while also considering significant contributions from Latin America that have methodological relevance. Table 1 summarizes the main references discussed.
2.1 Machine Learning in Labor Market Estimation
Supervised learning methods have increasingly been applied to estimate labor statistics in contexts with limited direct observation. These approaches utilize flexible algorithms and auxiliary covariates to infer labor outcomes in situations of incomplete coverage. A key advantage of these methods is their capacity to learn complex, nonlinear relationships from the data, enabling the models to generalize patterns that may not be evident using traditional econometric techniques.
van Dijk et al., (2022) developed high-resolution labor maps for Vietnam by downscaling district-level employment data using a super learner ensemble. Their model integrated satellite imagery, nightlight intensity, and household surveys to estimate spatial employment distributions across diverse areas. This study demonstrates the effectiveness of ensemble methods in interpolating labor variables in data-scarce contexts.
In Colombia, Vidal, Sierra-Suárez & Cerón, (2024) proposed a composite labor index that integrates informality rates, labor expectations, and behavioral indicators derived from Google Trends. Their model employs tree-based learners and feature selection techniques, effectively capturing short-term dynamics in urban labor markets despite the absence of high-frequency official data. Although their analysis was limited to city-level series, it demonstrates the predictive potential of nontraditional data sources.
Recent studies in Colombia have utilized supervised learning techniques to predict unemployment, labor force participation, and informality by incorporating economic variables, alternative data sources (e.g., Google Trends), and official survey data. Orozco-Castañeda et al., (2024) A composite labor indicator will be constructed through machine learning models trained on city-level inputs, highlighting the potential of these methods to capture complex urban labor dynamics. Pérez-Rosero et al., (2025) Explainability will be emphasized by integrating Uniform Manifold Approximation and Projection (UMAP) with Gaussian Processes to develop interpretable models of unemployment. These efforts underscore the increasing empirical relevance of machine learning techniques in Colombian labor research.
These studies confirm the feasibility of applying supervised learning to labor market modeling, provided that the models incorporate robust anchor variables and context-sensitive predictors.
2.2 Temporal Disaggregation for Labor Indicators
Temporal disaggregation techniques facilitate the estimation of high-frequency time series from low-frequency aggregates. These techniques are particularly pertinent when official labor statistics are available solely on an annual basis at the subnational level.
In contexts where low-frequency data are present, temporal disaggregation is employed to construct high-frequency labor series. In Colombia, González-Herrera, (2018) applied the Denton and Chow-Lin methods to interpolate employment data while maintaining consistency constraints. Although these methods are standard in national accounts, their application to social indicators is limited in practice. Sánchez & Morales, (2015) establishes the foundation for reconstructing intra-annual variation in labor aggregates when direct monthly data are unavailable.
Chow & Lin, (1971) introduced a regression-based approach that distributes annual data into monthly or quarterly values based on their covariation with an auxiliary indicator. Denton, (1971) proposed a quadratic minimization procedure that smooths the interpolated path while preserving the trend of the indicator. Fernández, (1981) generalized these techniques to accommodate integrated processes, and Litterman, (1983) introduced a stochastic extension incorporating autoregressive innovations.
Recent developments, such as Mosley et al., (2022), incorporate sparse regularization (LASSO) for the selection among multiple noisy predictors during the disaggregation process. This approach enhances robustness in high-dimensional settings, where collinearity and overfitting are prevalent risks.
However, traditional disaggregation methods typically assume complete data at the aggregate level and are infrequently applied at the subnational scale. Their adaptation to contexts with fragmented spatial and temporal coverage is constrained.
2.3 Labor Market Estimation under Subnational Data Gaps
A significant empirical challenge in Colombia is the lack of monthly labor data for departments not included in the GEIH. Existing studies have employed methods such as spatial extrapolation, synthetic estimation, and predictive modeling. For example, OIT & CEPAL, (2022) provides technical guidelines for constructing consistent subnational labor indicators using auxiliary data and partial survey coverage.These methods represent viable strategies for estimating labor indicators in regions lacking direct observation, particularly when aligned with annual benchmarks. These methods are effective for estimating labor indicators in regions without direct observation, especially when correlated with annual benchmarks.
2.4 Regional Labor Segmentation and Convergence
Understanding regional labor disparities necessitates modeling structural heterogeneity. García-Peña, (2017) analyzes convergence in informal employment across Colombian departments using spatial econometrics. Their findings indicate persistent segmentation and path dependence in subnational labor markets. Similar studies in Latin America have utilized clustering, principal components, or typology frameworks to categorize territories based on labor structures. These methods support region-specific modeling strategies and facilitate the identification of structural gaps in labor conditions across territories.
2.5 Accounting Identities in Predictive Labor Modeling
Labor indicators are interdependent through accounting identities (e.g., labor force = employed + unemployed). However, few empirical models explicitly enforce these constraints. Recent work, such as OECD, (2021), explores multi-output machine learning models and structural regularization to ensure coherence among predicted indicators. Incorporating such restrictions enhances interpretability and consistency, particularly when the outputs are utilized for policy formulation or official statistics. In Latin America, this area remains open for methodological development, particularly regarding the integration with machine learning pipelines.
Table 1: Selected literature on labor market estimation and machine learning (2010-2025)
Authors Year Country Main Contribution
Orozco-Castañeda et al., (2024) 2024 Colombia Forecasting unemployment using SVR and neural networks with economic indicators and Google Trends.
Vidal, Sierra-Suárez & Cerón, (2024) 2024 Colombia Construction of a regional labor market indicator via ML, combining official data with alternative sources in major cities.
Pérez-Rosero et al., (2025) 2025 Colombia Explainable ML approach (Gaussian Processes + UMAP) to predict unemployment dynamics with interpretability.
Fajardo et al., (2021) 2021 Colombia Prediction of business closures and employment risks using ensemble models (RF, XGBoost) on firm-level data.
Pizzinelli et al., (2023) 2023 LATAM Regional exposure to automation and AI via task-based occupational scoring across multiple countries.
González-Herrera, (2018) 2018 Colombia Temporal disaggregation of employment series using Denton and Chow-Lin under limited data settings.
Sánchez & Morales, (2015) 2015 Chile Spatiotemporal estimation of unemployment with incomplete coverage using Bayesian hierarchical models.
OIT & CEPAL, (2022) 2022 LATAM Guidelines for constructing comparable labor indicators with coverage gaps and disaggregation methods.
OECD, (2021) 2021 Global Survey of empirical strategies combining ML with labor economics and structural constraints in developing contexts.
García-Peña, (2017) 2017 Colombia Regional convergence of informality using spatial econometrics and departmental panel data.
van Dijk et al., (2022) 2022 Vietnam High-resolution labor mapping using super learner ensemble with satellite and survey data.
Note: Author’s own construction.
2.6 Contribution and Positioning of this Study
This study significantly contributes to the labor statistics literature by addressing longstanding territorial and temporal gaps in official measurements. First, it introduces a harmonized monthly panel of labor market indicators for all 33 Colombian departments, covering the period from 1993 to 2025. This representation effectively bridges a critical empirical gap in subnational labor monitoring. Second, it develops an integrated estimation pipeline that combines supervised machine learning, econometric disaggregation, deep learning, and accounting-based reconstruction. Third, it implements a structured calibration protocol, ensuring consistency with national aggregates while maintaining internal coherence across labor indicators.
3 Methodology
3.1 General overview of the multistage pipeline
Figure 1 summarizes the multi-stage pipeline developed to reconstruct monthly labor indicators for all Colombian departments from 1993 to 2025. This process is hierarchical, progressing from national anchors to city-level dynamics and finally to departmental series. At each stage, consistency is ensured through the integration of official data sources, demographic controls, and statistical learning methods.
National baselineWB/ILO + GEIH + demographic projectionsConcept harmonization and calibration to official aggregatesMonthly national disaggregationTemporal disaggregation to provide national anchorsMonthly reference series used to reconcile all subnational estimatesCity signals construction32 principal cities since 2021 + artificial CundinamarcaBackward splicing with 23 GEIH domains; correlation-based mappingCity monthly alignmentAnnual build then monthly interpolationCities act as predictive signals extended for 1993-2025Departmental reconstruction (23)Disaggregate annuals, enforce identities, demographic anchorsBackward extensions only when evidence availableML extrapolation to 33 deptsXGBoost with city dynamics + socioeconomic anchorsTrain: 23 depts; Predict: monthly 1993-2025 constrained to totalsInformality modelFeedforward NN on 23 cities 2007-2025Extrapolate to depts and reconcile with total employmentFinal consolidation and validationIdentities, alignment, smoothing, metricsEmployed + Unemployed = PEA; PEA + Inactive = PETValidate vs GEIH annuals and city benchmarks
Figure 1: Overview of the pipeline and key methods
Note: Author’s own construction.
3.2 National Labor Market Baseline
A consistent national baseline is established by compiling data from the World Bank, the International Labour Organization (ILO), the General Household Survey (GEIH, Gran Encuesta Integrada de Hogares), and official demographic projections. Concepts are harmonized, and totals are calibrated to official national aggregates. Monthly series are derived through temporal disaggregation, providing the reference against which all subnational reconstructions are reconciled. This process ensures that departmental estimates remain coherent with national labor accounts.
3.2.1 Adjustment of World Bank Labor Estimates
To extend coverage prior to the GEIH, we reconstruct Colombia’s national labor market series using international data from 1993 to 2000. Annual indicators from the World Bank’s World Development Indicators World Bank, (2025), which incorporate ILO estimates ILO, (2025), provide figures for employment, unemployment, and population. These indicators are temporally disaggregated and integrated with monthly GEIH data from 2001 onward, ensuring demographic consistency and continuity.
However, global estimation frameworks often fail to capture dynamics specific to Colombia. For example, the surge in unemployment following the 1998–1999 recession, the mild impact of the 2008-2009 financial crisis, and the sharp shock of COVID-19 in 2020 are only partially reflected Medina et al., (2013); Grosh et al., (2014); Alvarez & Pizzinelli, (2021); DANE, (2025); BanRep, (2025). To address these deficiencies, we structurally recalibrate the series by smoothing employment and unemployment to align with macroeconomic cycles, reconstructing participation based on its empirical relationship with employment, and deriving aggregates using official projections of the working-age population.
The resulting annual estimates encompass employment, unemployment, labor force (PEA), inactivity, and the working-age population (PET). All estimates are internally consistent, and their comparison with GEIH benchmarks (2001-2024) demonstrates that average absolute percentage errors (MAPE) are below 5%. These series function as macro anchors for monthly disaggregation, providing a harmonized baseline for the period 1993-2025.
3.2.2 Demographic Baseline and National Labor Indicators
The demographic baseline is based on official population projections published by DANE DANE, 2023a . These projections cover the period from 1993 to 2050 and are available in three vintages (1993-2004, 2005-2019, 2020-2050), which have been harmonized to ensure consistency. Departmental projections extend through 2035. These projections provide denominators for labor rates (e.g., participation and unemployment) and facilitate the conversion of rates into absolute levels that align with official totals.
In parallel, monthly labor indicators are compiled from the GEIH DANE, 2025a , producing a harmonized panel extending from 2001 to 2025. Together, the DANE projections and GEIH indicators establish the national trajectory that underpins all departmental estimates.
3.2.3 Monthly Disaggregation of World Bank Labor Data
To extend labor indicators beyond the GEIH window, annual World Bank estimates are disaggregated into a monthly frequency using macroeconomic covariates. Instead of interpolating absolute levels, we reconstruct ratios of employment, unemployment, participation (PEA), working-age population (PET), and inactivity relative to reference populations. This approach minimizes seasonal noise and inconsistencies among sources.
Monthly trajectories are derived using Chow–Lin models calibrated on macroeconomic predictors, including the Consumer Price Index (CPI, Índice de Precios al Consumidor), Producer Price Index (PPI, Índice de Precios al Productor), exchange rate (TRM, Tasa Representativa del Mercado), and real minimum wage. The selection of predictors is based on their empirical correlation with labor cycles Stock & Watson, (1999); Galí, (1999). The optimal predictor minimizes the Mean Absolute Percentage Error (MAPE) relative to GEIH benchmarks from 2001 to 2024, with all ratios resulting in errors below 1.5.
Ratios are converted into levels using DANE projections. PET is linearly smoothed to avoid discontinuities, while other aggregates are reconstructed utilizing accounting identities. Coherence is maintained consistently, ensuring that the sum of employed and unemployed individuals equals PEA, and the sum of PEA and inactive individuals equals PET. Figures B.1 and B.3 illustrate the reconstructed ratios and residual adjustments. The result is a demographically consistent, high-frequency national baseline for the years 1993-2025, which serves as the foundation for all departmental reconstructions. Comprehensive technical details, including equations, diagnostics, and sensitivity checks, are presented in Appendix B.
3.3 Labor Market Data Reconstruction and Preparation
3.3.1 City-level Implementation
The reconstruction process of the subnational labor market begins with the 32 principal cities, for which monthly labor market indicators have been available since December 2021, alongside a synthetic series for Cundinamarca (excluding Bogotá). The signals from these cities are extended backward using the 23 GEIH urban domains as references. In instances where a city lacks a direct match, the backward extension employs the domain exhibiting the highest correlation. To mitigate noise, the extensions are initially calculated at an annual frequency and subsequently disaggregated into monthly series that align with national totals. This approach results in a coherent framework of 33 city-level signals, continuously extended from 1993 to 2025, which serve as predictors for the departmental reconstruction.
City-department mappings are established based on pairwise correlations of monthly indicators. Spearman’s rank correlation serves as the primary criterion, supplemented by Kendall’s tau as a robustness check. Pairs with fewer than ten overlapping observations are excluded from the analysis. For departments lacking direct observations, the donor with the strongest correlation is selected. In the case of Cundinamarca, departmental results are utilized directly to ensure comprehensive coverage. This approach results in a stable donor structure applicable across all extensions.
The 32 observed cities and the synthetic Cundinamarca series are aggregated to annual frequency (medians, 2021-2025) and linked to GEIH-compatible domains. Direct matches are paired directly, while unmatched cities utilize donor series. Proportional retropolarization extends these annual signals back to 2007, ensuring consistent levels and growth rates. Monthly disaggregation is then performed using the optimal Chow–Lin method (average conversion), anchored to national monthly indicators. The final panel for 1993-2025 is produced through proportional splicing against national totals, ensuring consistency in population, employment, unemployment, inactivity, PEA, and PET. Labor accounting identities are enforced, and national projections are aligned. Full implementation details are provided in  Appendix C.
3.3.2 Departmental Implementation
The second stage addresses the departments, which are the primary targets for estimation. For the 23 departments directly covered by the GEIH, official annual aggregates are disaggregated, labor accounting identities are applied, and demographic projections are integrated. For the 23 departments covered by the GEIH (2007-2024), annual aggregates are converted to a monthly frequency using the Chow-Lin method, which is anchored to national indicators. This approach preserves both annual totals and temporal continuity.
Historical extensions to 1993 are included only when sufficient evidence is available (threshold of 0.95), thus avoiding artificial interpolations. Unlike city-level predictors, which are systematically extended to provide inputs for modeling, departmental series are treated as final outputs and, therefore, necessitate higher accuracy. These reconstructed series serve as the dependent variables in the predictive stage, with city signals and other covariates functioning as explanatory inputs.
3.4 Data Integration
3.4.1 Covariate Construction
A unified set of monthly predictors is compiled to model departmental labor indicators through supervised learning. Covariates are selected based on three criteria: (i) availability since 1993, (ii) monthly frequency, and (iii) empirical relevance to labor market dynamics.
Macroeconomic predictors include the nominal exchange rate (TRM, Tasa Representativa del Mercado), consumer price index (CPI, Índice de Precios al Consumidor), producer price index (PPI, Índice de Precios al Productor), exports, imports, real value unit (UVR, Unidad de Valor Real), and the industrial production index (IPI, Índice de Producción Industrial). Institutional predictors consist of the legal minimum wage (Salario Mínimo Legal Vigente) and the transport subsidy (Subsidio de Transporte). A real minimum wage proxy is derived by deflating the nominal wage using the CPI. GDP and departmental production are excluded due to their annual frequency and delayed update. Consequently, the final set of covariates primarily comprises monthly indicators, supplemented by a limited number of retrospectively expanded annual series. Comprehensive definitions and data sources are reported in Appendix E.
3.4.2 Cluster-Based Structural Variations
To enhance generalization, departments are grouped into clusters that capture structural similarities in demographic, geographic, and economic characteristics. Clustering allows unobserved departments to borrow information from observed departments, thereby incorporating variability that is not entirely explained by labor indicators . A strict constraint is imposed: every cluster must contain at least one department with GEIH coverage. This constraint prevents the formation of synthetic clusters composed exclusively of unobserved units, ensuring that all extrapolations are empirically grounded (Table E.2).
3.4.3 Preparation of Labor Series
Model outputs are initially derived as monthly labor shares (employment, unemployment, and inactivity). These shares are converted into absolute population counts using official demographic projections and accounting identities: employment = employment rate $\times$ PET, unemployment = unemployment rate $\times$ PEA, and inactivity = inactivity rate $\times$ PET. This approach ensures consistency between rates and levels over time.
For departments covered by the GEIH, splicing aligns the observed and reconstructed series prior to level conversion. For unobserved departments, the trained model generates monthly shares, which are transformed in a similar manner. Details regarding the splicing formulation and reconstruction logic are provided in Appendix F.
3.4.4 Data Consolidation
All elements are integrated into a department-month feature matrix suitable for supervised estimation. Two partitions are defined: (i) a training set consisting of 23 departments with observed labor outcomes and (ii) a prediction set comprising 33 departments, which includes both observed and unobserved cases. This setup facilitates out-of-sample predictions, with generalization validated through cross-department predictive checks. The variables were selected by prioritizing those with high frequency and timely updates, ensuring that the methodological framework allows for the continuous and up-to-date reconstruction of indicators.
3.5 Model for Main Labor Market Indicators
3.5.1 Prediction of Departmental Labor Indicators (1993-2025)
To reconstruct monthly labor indicators for the 33 departments from 1993 to 2025, we employ a supervised learning framework for multi-output prediction. This framework captures non-linear dynamics, leverages shared structures across outputs, and integrates heterogeneous predictors. The five core targets, working-age population (PET, Población en Edad de Trabajar), economically active population (PEA, Población Económicamente Activa), employed individuals, unemployed individuals, and inactive individuals, are estimated jointly, ensuring that temporal and structural patterns are consistently modeled. The framework relies on the strength of city-level signals, which provide sufficiently rich information to extrapolate departmental trajectories without necessitating more complex architectures.
The modeling strategy employs gradient-boosted regression trees (XGBoost). Each indicator is formulated as a function of demographic and institutional covariates. Targets are modeled as population shares to enhance stability and comparability across departments. Following estimation, absolute levels are derived by rescaling predicted shares with the population feature, which is intentionally excluded from inputs to prevent data leakage. The boosting process iteratively fits residuals, enabling the model to capture complex interactions and non-linearities with high predictive accuracy and generalization capabilities.
Training follows systematic procedures, including feature scaling to stabilize optimization. Performance is evaluated using three standard metrics: Mean Absolute Error (MAE), Mean Absolute Percentage Error (MAPE), and Root Mean Squared Error (RMSE). These metrics provide a comprehensive assessment of accuracy and enable comparability across indicators and departments.
Target transformation and scaling. Targets are normalized as shares to enhance numerical stability and comparability across departments and over time, while preventing data leakage from absolute levels. After making predictions, absolute counts are recovered by rescaling the estimated shares using official population projections for the period 1993-2035.
Input features. The design matrix incorporates demographic, macroeconomic, and labor-related signals:
  • Demographic inputs: Official departmental population projections shares (urban and rural), by year.
  • Macroeconomic variables: Monthly national indicators including exchange rate (TRM, Tasa Representativa del Mercado), consumer price index (CPI, Índice de Precios al Consumidor), producer price index (PPI, Índice de Precios al Productor), inflation, real value unit (UVR, Unidad de Valor Real), imports, exports, and real minimum wage (Salario Mínimo Legal Vigente deflated by CPI).
  • Labor signals: Monthly labor indicators from GEIH (Gran Encuesta Integrada de Hogares), available only for directly observed departments.
  • Structural clusters: Cluster-level averages of labor indicators, acting as proxy dynamics for departments without GEIH data.
Training and validation strategy. All features and targets are standardized prior to training. The training set encompasses the 23 GEIH departments. Model validation adheres to four schemes:
  1. 1.
    No-holdout fit: Full-sample training, providing an upper bound of fit.
  2. 2.
    Stratified holdout: 80-20 split by year, testing temporal generalization.
  3. 3.
    Leave-$k$-out (LKO): Random exclusion of $k=3$ departments, testing robustness under partial spatial exclusion.
  4. 4.
    Leave-one-group-out (LOGO): Iterative exclusion of each department, simulating extrapolation to unobserved units.
Performance is evaluated based on share-based targets. The stratified holdout method yields an average Mean Absolute Percentage Error (MAPE) of 1.38%. LOGO errors increase to 10–12%, highlighting the challenges associated with spatial extrapolation in the absence of direct anchors. These results are accessible on G.3.
Prediction and generalization. The final model is trained on all observed departments and generates monthly predictions for all 33 departments. For departments that lack GEIH data, results are primarily influenced by macroeconomic predictors, demographic projections, and cluster-level proxies.
Post-processing and alignment. Predictions are adjusted to ensure demographic and national consistency:
  1. 1.
    National alignment: Departmental shares are proportionally scaled to match reconstructed national totals.
  2. 2.
    Conversion to levels: Rescaled shares are multiplied by departmental population projections, yielding absolute counts consistent with labor accounting identities.
The resulting panel yields internally consistent departmental labor indicators that align with national benchmarks, rendering it suitable for longitudinal analysis and policy applications. All technical details can be found on Appendix G, which includes a description of the estimation procedure, an evaluation of model performance, relevant metrics, and a visual comparison of the primary labor market results against official annual data.
3.6 Informality Rate Estimation with Neural Networks
Informality presents a unique challenge due to the absence of official departmental series to serve as benchmarks. While employment, unemployment, and inactivity can be reconstructed by mapping city-level survey data to departments, informality is reported only for 23 metropolitan areas in the GEIH (Gran Encuesta Integrada de Hogares), with no official departmental totals. This lack of data necessitates the extrapolation of urban evidence to departments with significant rural sectors, where the dynamics of informality differ substantially. Consequently, this situation requires a more complex approach than a typical optimized machine learning model.
To address this gap, we implement a custom neural network framework specifically designed to capture the heterogeneous and non-linear drivers of informality. Standard methods, such as gradient boosting or baseline multilayer perceptrons, tend to regress toward the mean and perform poorly in regions of high informality (MAPE $18\%$ on average). Our architecture introduces domain-informed inductive biases and robust training strategies, thereby enhancing extrapolation to departments lacking direct survey coverage.
3.6.1 National Informality Reconstruction
A consistent national benchmark for the years 1993 to 2025 is established through four steps:
  1. 1.
    City-level harmonization: Informality rates for the 23 cities are harmonized with spline-based interpolation (PCHIP and Akima) and seasonal adjustment.
  2. 2.
    National series backcasting: Informality counts are spliced with reconstructed employment totals using stable regressions, producing a continuous national trajectory from 1993 onward.
  3. 3.
    City-level backcasting: City series are extended backwards in alignment with the national benchmark, avoiding artificial volatility.
  4. 4.
    Neural network calibration: These harmonized city and national benchmarks provide the training and calibration targets for the departmental model.
This process enforces coherence across different levels, prevents spurious breaks, and ensures additivity to the national employment base.
3.6.2 Neural Network Architecture and Training
The departmental model is implemented as a custom multilayer perceptron (MLP) that incorporates residual connections, dropout, and optional monotonicity constraints. Key elements include:
  • Residual blocks: Improve gradient flow and stabilize training over long historical series.
  • Monotonic layers: Encode prior knowledge by constraining weights of variables such as education or minimum wage to monotonic effects.
  • Custom loss functions: Networks are trained either on logit-transformed rates with Huber or MSE loss, or directly in probability space with a Beta likelihood, improving performance near boundary values.
  • Group-aware validation: Cross-validation protocols (temporal 80/20, Leave-One-Group-Out, Leave-$k$-Out, and Leave-One-Year-Out) are built into the pipeline to test robustness under both spatial and temporal extrapolation.
Training employs early stopping, adaptive learning rates, and systematic input scaling. Predictions are limited to $[0,1]$, ensuring interpretability as rates.
3.6.3 Post-processing and Validation
Predicted departmental informality rates are converted into counts of informal workers by multiplying them with reconstructed employment levels. Monthly rescaling aligns departmental totals with the national benchmark, after which the rates are reinstated. This process ensures additivity and consistency across levels.
Validation is conducted using four protocols: (i) in-sample fit, (ii) temporal 80/20 split, (iii) Leave-One-Group-Out (LOGO), and (iv) Leave-$k$-Out (LKO). The custom neural network consistently outperforms simpler methods, particularly in high-informality regimes where baseline models exhibit failure. Results are summarized in Table H.4.
In conclusion, the neural network framework offers nationally consistent estimates of departmental informality. Comprehensive details, including the formal specification, MLP architecture, optimization strategies, calibration and validation protocols, and visual comparisons with official monthly data, are presented in Appendix H.
3.6.4 Labor Market Profiles and Clustering
To assess departmental labor market performance, we construct a composite Employment Quality Index (EQI). The EQI consolidates multiple indicators into a standardized monthly metric that reflects both the level and structural quality of employment. A penalty adjustment is included to account for scenarios in which favorable outcomes, such as high employment rates, coexist with a significant degree of informality.
Each underlying indicator is smoothed to reduce seasonality and to emphasize long-term dynamics. It is then converted into percentile scores, with directionality defined by whether higher values indicate improvements (e.g., employment, participation) or deteriorations (e.g., unemployment, informality). Standardized indicators are aggregated using normalized weights, which are applied uniformly unless specified otherwise. Subsequently, departments are classified into ordered categories—Very Low, Low, Medium, High, and Very High—based on their average Employment Quality Index (EQI), resulting in a typology of labor market quality across regions.
To analyze dynamics, we compute periodic averages and departmental rankings, which are visualized through bump charts that illustrate improvement, stagnation, or decline over time. In addition to ranking, the EQI supports the segmentation of departments by clustering regions based on long-run average EQI, thereby creating comparable profiles of employment quality. This framework is used to contextualize trajectories and facilitate peer comparisons (see Table 8). Comprehensive methodological details, including indicator selection, weighting, smoothing procedures, and robustness checks, are provided in Appendix I.
3.7 Methodological Summary and Transition to Analysis
The methodological framework is structured in seven interconnected stages: establishing a national baseline with international benchmarks and GEIH information; reconstructing city-level signals and extending them with historical domains; generating departmental series with demographic consistency and accounting identities; compiling a unified feature set of macroeconomic and institutional covariates; predicting departmental labor indicators with a supervised XGBoost model; estimating informality through a tailored neural network with residual connections and monotonicity constraints; and constructing the Employment Quality Index (EQI) to integrate level and quality dimensions of labor outcomes.
This dataset facilitates a systematic examination of national and subnational dynamics, encompassing patterns of convergence and divergence, as well as heterogeneous responses to significant shocks, such as the 1999 financial crisis, the 2008 global recession, and the COVID-19 pandemic. By employing methodologies such as trend decomposition, correlation analysis, Granger causality tests, composite index construction, and comparative visualization, the dataset elucidates the interplay among participation, employment, unemployment, inactivity, and informality over time and across regions. Beyond its descriptive capacity, the dataset establishes an empirical foundation for policy by highlighting persistent gaps in employment quality, identifying clusters of departments with shared trajectories, and supporting targeted interventions to reduce informality, enhance participation, and strengthen territorial resilience.
4 Results and Discussion
The reconstruction pipeline produces a harmonized and internally consistent set of labor market indicators that ensure comprehensive temporal and spatial coverage. The final outputs comprise:
  • National monthly series (1993-2025): one time series per indicator, consistent with official aggregates where available.
  • Departmental monthly series (1993-2024): estimates for all 33 departments, covering observed and reconstructed periods.
  • Annual departmental aggregates: computed as the sum of the 12 monthly estimates for each year and department.
  • Annual national aggregates: computed as the cross-departmental sum of monthly estimates with exact aggregation.
These outputs provide a statistical foundation for labor analysis, territorial diagnostics, and the design of employment policies. Depending on the analytical horizon, we recommend three standard panels:
  1. 1.
    Full historical panel (1993-2025): long-run structure, historical cycles, convergence.
  2. 2.
    Post-2001 panel: aligned with GEIH microdata, medium-term coverage with higher reliability.
  3. 3.
    High-reliability panel (2007-2025): post-redesign GEIH period with maximum cross-department comparability.
A design principle applicable to all products is demographic and accounting consistency. Official population projections serve as
  • Denominators to compute rates from counts.
  • Scaling anchors to recover absolute levels from predicted rates.
This approach preserves additivity across both temporal and geographic dimensions while maintaining internal identities.
Unlike employment, unemployment, or inactivity, the informality rate lacks an official departmental counterpart. Consequently, departmental data are entirely extrapolated by the neural network. These series represent statistically consistent reconstructions rather than direct survey values.
4.1 Results of the Trained Models
4.1.1 Main Labor Market Series
Following the complete estimation and calibration pipeline, we evaluated accuracy by comparing the reconstructed annual departmental aggregates against the official GEIH results. This assessment captures the definitive reconstruction error, reflecting both the supervised predictions and the post-processing adjustments, which include the enforcement of accounting identities and national calibration.
Table 2 presents the Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE) for each indicator. These metrics are expressed in absolute counts and averaged across all departments and years from 2007 to 2024.
The results indicate strong predictive accuracy. The overall Mean Absolute Percentage Error (MAPE) across all variables is 1.8%. Errors are lowest for the employed and economically active population (PEA), both below 1.3%. Slightly higher errors are observed in the inactive population (PET, 1.5%) and the total population (2.3%), reflecting the compounding effect of demographic inputs. Unemployment exhibits a MAPE of 2.2%, consistent with its smaller scale and greater volatility.
Table 3 disaggregates MAPE by department and indicator. Performance is consistent across regions: in most departments, errors for core indicators remain below 3%, with employed, PEA, and PET typically ranging between 1.0% and 1.6%. The largest deviations are observed in Cundinamarca (up to 4.4%) and Quindío (2.9%), although these remain within acceptable thresholds. Overall, the pipeline demonstrates reliable spatial generalization and robust accuracy in subnational labor market estimation.
Performance metrics are computed solely for the 23 departments covered by the GEIH. For the remaining 10 departments, the lack of ground-truth values precludes direct error measurement. As a result, evaluation for these regions depends on structural plausibility, consistency with national totals, and the generalization properties of the model.
MAPE is emphasized as the primary metric because it provides scale-invariant comparisons across indicators and territories. Absolute deviations (e.g., $\pm$10,000 individuals) are less informative due to the heterogeneity in departmental population sizes, while percentage-based errors facilitate more interpretable assessments of relative accuracy.
Table 2: Validation Metrics by Variable (2007-2024).
Variable RMSE MAE MAPE (%)
Employed 15,975.2 9,487.5 1.3
Inactive 15,654.4 9,717.2 2.2
Unemployed 4,953.3 2,164.8 2.2
PEA 16,999.2 9,709.7 1.2
PET 31,047.6 18,941.5 1.5
Population 60,404.2 40,974.0 2.3
Total 24,172.3 15,165.8 1.8
Note: Author’s calculations.
Table 3: MAPE (%) by Department and Labor Market Variable (2007-2024).
Department Employed Inactive Unemployed PEA PET Population
Antioquia 1.2 2.4 2.5 1.3 1.7 2.8
Atlántico 1.1 2.2 1.6 1.2 1.6 2.7
Bolívar 1.0 2.1 2.4 1.0 1.4 2.5
Boyacá 1.4 1.8 1.6 1.2 1.4 1.9
Caldas 1.1 2.1 2.4 1.0 1.4 2.5
Caquetá 1.0 2.1 2.5 0.7 1.2 2.3
Cauca 1.2 1.7 1.6 1.0 1.2 1.8
Cesar 1.1 1.9 1.5 0.9 1.2 2.2
Córdoba 1.3 1.9 1.9 1.0 1.1 1.9
Cundinamarca 2.6 4.4 3.3 2.5 3.0 4.1
Chocó 2.2 2.9 2.4 2.1 2.3 2.6
Huila 1.5 1.7 2.2 1.2 1.4 1.9
La Guajira 1.1 2.6 2.6 1.1 1.3 2.0
Magdalena 1.0 1.8 1.3 0.9 1.2 2.1
Meta 1.6 2.3 2.3 1.3 1.6 2.0
Nariño 1.9 2.2 3.6 1.6 1.8 2.1
Norte de Santander 1.2 1.7 2.1 1.1 1.3 2.0
Quindío 1.7 2.5 2.9 1.4 1.8 2.9
Risaralda 1.3 2.4 2.1 1.1 1.6 2.6
Santander 0.9 2.0 1.8 0.8 1.2 2.3
Sucre 1.1 1.7 1.5 1.0 1.1 2.1
Tolima 1.4 2.1 2.7 0.9 1.2 2.0
Valle del Cauca 1.0 2.0 2.4 0.8 1.2 2.2
Note: Author’s calculations.
The distribution of estimation errors is uniform across departments, highlighting the model’s ability to generalize across heterogeneous regional labor markets. Most departments demonstrate Mean Absolute Percentage Errors (MAPEs) below 2.5% for the majority of indicators, with only a few instances exceeding 3%. This uniformity is significant considering the structural heterogeneity in demographic size, economic base, and labor informality across Colombian regions.
The estimation pipeline shows strong spatial adaptability. Predictive accuracy is preserved in both large and diverse departments, such as Antioquia, Santander, and Valle del Cauca, as well as in regions with more unstable dynamics, such as Chocó and Meta. This suggests that the model effectively captures both national and subnational structural patterns without overfitting to specific territories.
These results provide empirical validation of the modeling strategy. The low and stable MAPE values across departments and indicators suggest that the model captures structural relationships rather than merely reproducing noise or idiosyncratic fluctuations. The consistency of errors, even in departments with contrasting economic conditions, demonstrates the adaptability of the pipeline and its ability to generalize beyond the observed domains while utilizing all available data.
A detailed graphical comparison between annual observed and estimated values for each indicator is included in  Appendix G.5 These plots confirm the alignment of the reconstructed series with GEIH aggregates across departments. The coherence of levels and trends supports the reliability of the estimation pipeline, following calibration and enforcement of identity. Each figure contrasts annual predictions with observed values, providing visual evidence of the performance.
4.1.2 Informality Estimation Results
To estimate monthly labor informality at the departmental level, we first reconstructed a continuous time series and benchmarked it against observed data from the 23 primary cities for the period 2007-2025. This benchmarking enables a direct evaluation of the model’s ability to replicate observed labor market dynamics within the sample.
The in-sample performance across all cities indicates high accuracy. The RMSE of 0.0149 implies that deviations between predicted and observed informality rates average 1.49 percentage points on the original scale. The mean MAPE of 2% confirms that relative errors consistently remain low compared to the magnitude of the observed rates.
At the city level, MAPE values exhibit a narrow range around the 2% band. The smallest errors are observed in Santa Marta (1.40%), Florencia (1.66%), and Bucaramanga A.M., while the median across cities is 1.96%. Most results cluster near this median value, indicating a strong overall fit; however, a few cities, such as Bogotá D.C. and Barranquilla A.M., demonstrate slightly higher deviations that suggest potential for targeted refinements.
As in other labor market evaluation results, error metrics are computed solely for the 23 primary cities covered by the GEIH, which are the only domains with consistently high-frequency ground-truth data. For departments outside this coverage, direct error quantification is not feasible. In such cases, validation relies on indirect checks of structural plausibility, consistency with national and regional aggregates, and the expected generalization capacity inferred from the city-based sample.
The convergence of findings across these schemes supports the robustness of the estimation pipeline and its applicability in reconstructing monthly informality and labor market series beyond the directly observed domains. A visual comparison of observed and reconstructed informality rates across the 23 cities is presented in Figure H.11, thereby reinforcing the previously discussed quantitative results.
Table 4: Overall in-sample performance metrics for monthly informality estimation.
Scope RMSE MAE MAPE (%)
Informality Rate (%) 0.0149 0.0111 2.0045
Note: Author’s calculations.
Table 5: In-sample RMSE, MAE and MAPE (%) per main city and metropolitan areas (2007-2024).
City / Metro Area RMSE MAE MAPE
Armenia 0.015 0.011 2.32
Barranquilla A.M. 0.019 0.015 2.58
Bogotá D.C. 0.014 0.012 3.11
Bucaramanga A.M. 0.011 0.008 1.68
Cali A.M. 0.011 0.009 1.70
Cartagena 0.016 0.012 1.96
Cúcuta A.M. 0.016 0.013 2.01
Florencia 0.013 0.011 1.66
Ibagué 0.013 0.010 1.77
Manizales A.M. 0.011 0.009 2.24
Medellín A.M. 0.012 0.010 2.25
Montería 0.014 0.012 1.74
Neiva 0.016 0.013 2.28
Pasto 0.014 0.011 1.80
Pereira A.M. 0.013 0.009 1.98
Popayán 0.017 0.012 2.02
Quibdó 0.020 0.014 2.17
Riohacha 0.017 0.012 1.81
Santa Marta 0.009 0.007 1.04
Sincelejo 0.021 0.013 1.88
Tunja 0.016 0.011 2.41
Valledupar 0.017 0.013 1.89
Villavicencio 0.013 0.011 1.80
Note: Author’s calculations.
4.1.3 Synthesis of Validation Results
Table 2 reports the validation metrics for the six core labor variables reconstructed through temporal disaggregation and multi-output regression. For these variables, the Mean Absolute Percentage Error (MAPE) remains below 2.3%, with the lowest relative errors observed for Participation in the Economically Active Population (1.2%) and employment (1.3%).The aggregate error across all variables results in a MAPE of 1.8%.
For the informality rate, estimated using the neural network approach, the in-sample evaluation over the same period yields an RMSE of 0.0149 (in proportion units), an MAE of 0.0111, and a MAPE of 2% (Table 4). These values are directly comparable to the relative error magnitudes obtained for the reconstructed variables, thereby reinforcing the internal consistency of the modeling framework.
The in-sample results and the complementary validation schemes provide strong and convergent evidence for the stability and generalization capacity of the proposed estimation framework. The consistency of error magnitudes across cities, variables, and test protocols indicates that predictive performance is not an artifact of a specific calibration subset but a consequence of structurally robust modeling choices.
4.2 National Labor Trends (1993-2025)
This subsection presents monthly national estimates of key labor market aggregates and rates from 1993 to 2025. The series are obtained by aggregating the reconstructed departmental indicators and serve as the benchmark reference for the estimation framework. They integrate official population projections, GEIH microdata, and historical benchmarks from the World Bank and ILO, harmonized through supervised learning, temporal disaggregation, and accounting-consistent smoothing techniques.
Figure 2 displays six national aggregates: working-age population (PET), total population, economically active population (PEA), employed individuals, unemployed individuals, and inactive individuals. PET and total population exhibit stable long-term growth, consistent with demographic projections. Employment and inactivity increase gradually, reflecting both the expansion of the labor market and evolving patterns of participation. Informal employment declines steadily until the late 2010s, while formal employment experiences growth.
The COVID-19 pandemic disrupted these trends. Both informal and formal employment fell sharply, followed by a quicker recovery of informal jobs, which is consistent with their greater flexibility and lower entry barriers. The PEA and employment series reflect these disruptions: a contraction in 2020, a surge in inactivity, and temporary exits from the labor force. Unemployment reached historically high levels before stabilizing around pre-pandemic values by 2022. Informality, which had been declining, rose temporarily during the recovery, highlighting the role of informal job creation in the economic rebound.
Figure 4 presents the corresponding national rates. The participation rate increases from the mid-1990s to the early 2000s, stabilizing near 65% thereafter. This trend is consistent with higher education levels, increased female participation, and urbanization. The unemployment rate fluctuates between 9% and 15%, peaking during the 1999 financial crisis and the 2020 pandemic, followed by gradual recoveries. The informality rate declines from above 68% in the early 1990s to approximately 56% by 2024, exhibiting cyclical fluctuations and a rebound related to the pandemic. The inactivity rate mirrors participation and remains above 35%, highlighting persistent barriers to inclusion.
Overall, the national indicators demonstrate credible long-term dynamics, coherent responses to shocks, and consistency with demographic constraints and accounting identities. Informality has exhibited a sustained decline over three decades, interrupted only by significant crises, particularly the COVID-19 pandemic, after which the downward trend resumes. These patterns confirm the structural coherence of the reconstructed series and suggest persistent regional heterogeneity.
Figure 2: National labor market levels (1993-2025): monthly estimates of working-age population and total population.
Note: Author’s calculations.
Figure 3: National labor market levels (1993-2025): monthly estimates of labor force, employment, unemployment, formals, informals and inactives.
Note: Author’s calculations.
Figure 4: National labor market rates (1993-2025): participation, employment, unemployment, informality, and inactivity rates.
Note: Author’s calculations.
Table 6: Monthly National Labor Market Statistics (1993-2025).
Variable Min Max Mean Median Std CV
employment_rate 0.427 0.660 0.587 0.587 0.032 0.055
unemployment_rate 0.071 0.216 0.120 0.114 0.028 0.235
participation_rate 0.534 0.719 0.667 0.668 0.027 0.042
inactivity_rate 0.281 0.466 0.333 0.332 0.028 0.083
informality_rate 0.553 0.708 0.627 0.628 0.043 0.069
Note: Author’s calculations.
The table 6 summarizes national monthly labor market statistics for the period 1993-2025. The employment rate ranges from 0.427 to 0.660, with a mean of 0.587 and a low coefficient of variation (CV) of 5.5%, indicating stable labor absorption over the long term. The unemployment rate spans from 0.071 to 0.216, averaging 0.120, but exhibits the highest variability (CV = 23.5%), confirming its sensitivity to cyclical shocks. The participation rate lies between 0.534 and 0.719, with a mean of 0.667 and minimal variation (CV = 4.2%). Inactivity ranges from 0.281 to 0.466, averaging 0.333 (CV = 8.3%), as expected due to its complementarity with participation. Informality fluctuates between 0.553 and 0.708, averaging 0.627, with low dispersion (CV = 6.9%). Overall, participation, employment, and informality remain structurally stable, while unemployment is the most volatile indicator.
Table 7 compares the response of labor indicators to four major macroeconomic shocks occurring between 1993 and 2025, using the two years prior to each event as a baseline. This approach captures both the magnitude and direction of short-term disruptions, as well as structural shifts.
The 1999 financial crisis is notable for being the most significant contraction. In comparison to 1996-1997, participation and employment, accompanied by a sharp rise in unemployment and inactivity. In addition to the job losses associated with the banking collapse, voluntary withdrawals from the labor force exacerbated the downturn. Although informality increased only modestly, its already high baseline indicates a structural deterioration in job quality. The prolonged recovery further underscores the severity of the shock.
The 2008 global recession resulted in only mild adjustments. In comparison to the years 2006-2007, participation and employment increased slightly, unemployment remained stable, and inactivity decreased marginally. Informality also decreased, consistent with favorable external conditions and counter-cyclical policies. No evidence of long-term scarring has been observed.
The COVID-19 pandemic resulted in an abrupt and synchronized economic collapse. In comparison to the years 2018-2019, participation and employment declined sharply, unemployment surged, and inactivity increased. Informal employment rose after a decade of decline, highlighting the vulnerability of urban service-sector jobs to lockdowns and the limited capacity for telework. This episode combined immediate disruption with regressive structural effects.
The post-pandemic recovery (2021-2022 baseline) indicates partial normalization. Participation and employment rates have improved, unemployment has decreased and inactivity has declined. Informality has returned to pre-pandemic levels; however, persistently high rates in several metropolitan areas underscore the concentration of new jobs in lower-quality segments.
Taken together, these episodes reveal distinct dynamics: the 1999 crisis caused long-term scarring, the 2008 recession was absorbed with limited adjustment, the pandemic exposed systemic vulnerabilities, and the recovery underscored the gap between quantitative rebounds and qualitative improvements. These contrasts suggest the need for differentiated policy strategies: structural reconstruction in deep crises, countercyclical stabilization in moderate recessions, and hybrid approaches that combine emergency relief with long-term reform in response to systemic shocks.
Table 7: Major Labor Market Shocks in Colombia (1993-2025): rows list metric-phase blocks; columns are shocks. Values are multi-year window averages and deltas vs the pre-shock baseline. For 2025, the ’after’ phase uses data available.
Shock 1999 Financial Crisis 2008 Global Recession COVID-19 Pandemic Post-COVID Recovery
Metric / Phase
Participation Rate (Before) $70.9\%$ $64.1\%$ $65.2\%$ $62.6\%$
Participation Rate (During) $67.8\%$ $65.2\%$ $\% 64.064.0\%$
Participation Rate - $\Delta$ During-Before $\downarrow\;-3.2\,\mathrm{pp}$ $\uparrow\;+1.1\,\mathrm{pp}$ $\downarrow\;-4.8\,\mathrm{pp}$ $\uparrow\;+1.4\,\mathrm{pp}$
Participation Rate (After) $67.5\%$ $67.9\%$ $62.6\%$ $64.2\%$
Participation Rate - $\Delta$ After-Before $\downarrow\;-3.4\,\mathrm{pp}$ $\uparrow\;+3.8\,\mathrm{pp}$ $\downarrow\;-2.6\,\mathrm{pp}$ $\uparrow\;+1.6\,\mathrm{pp}$
Employment Rate (Before) $61.5\%$ $56.6\%$ $\% 54.854.8\%$
Employment Rate (During) $56.7\%$ $57.5\%$ $50.5\%$ $57.5\%$
Employment Rate - $\Delta$ During-Before $\downarrow\;-4.8\,\mathrm{pp}$ $\uparrow\;+0.9\,\mathrm{pp}$ $\downarrow\;-7.9\,\mathrm{pp}$ $\uparrow\;+2.7\,\mathrm{pp}$
Employment Rate (After) $57.7\%$ $60.1\%$ $54.8\%$ $58.0\%$
Employment Rate - $\Delta$ After-Before $\downarrow\;-3.7\,\mathrm{pp}$ $\uparrow\;+3.5\,\mathrm{pp}$ $\downarrow\;-3.6\,\mathrm{pp}$ $\uparrow\;+3.2\,\mathrm{pp}$
Unemployment Rate (Before) $3.4\% 11$$1.7\%$ $10.4\%$ $12.5\%$
Unemployment Rate (During) $6.3\% 11$$1.8\%$ $16.6\%$ $10.2\%$
Unemployment Rate - $\Delta$ During-Before $\uparrow\;+3.0\,\mathrm{pp}$ $\leftrightarrow\;+0.1\,\mathrm{pp}$ $\uparrow\;+6.1\,\mathrm{pp}$ $\downarrow\;-2.3\,\mathrm{pp}$
Unemployment Rate (After) $4.4\% 11$$1.5\%$ $12.5\%$ $9.7\%$
Unemployment Rate - $\Delta$ After-Before $\uparrow\;+1.1\,\mathrm{pp}$ $\leftrightarrow\;-0.1\,\mathrm{pp}$ $\uparrow\;+2.1\,\mathrm{pp}$ $\downarrow\;-2.8\,\mathrm{pp}$
Inactivity Rate (Before) $29.1\%$ $35.9\%$ $34.8\%$ $37.4\%$
Inactivity Rate (During) $32.2\%$ $34.8\%$ $39.6\%$ $36.0\%$
Inactivity Rate - $\Delta$ During-Before $\uparrow\;+3.2\,\mathrm{pp}$ $\downarrow\;-1.1\,\mathrm{pp}$ $\uparrow\;+4.8\,\mathrm{pp}$ $\downarrow\;-1.4\,\mathrm{pp}$
Inactivity Rate (After) $32.5\%$ $32.1\%$ $37.4\%$ $35.8\%$
Inactivity Rate - $\Delta$ After-Before $\uparrow\;+3.4\,\mathrm{pp}$ $\downarrow\;-3.8\,\mathrm{pp}$ $\uparrow\;+2.6\,\mathrm{pp}$ $\downarrow\;-1.6\,\mathrm{pp}$
Informality Rate (Before) $67.6\%$ $64.3\%$ $58.1\%$ $58.5\%$
Informality Rate (During) $68.1\%$ $62.9\%$ $60.7\%$ $56.1\%$
Informality Rate - $\Delta$ During-Before $\uparrow\;+0.5\,\mathrm{pp}$ $\downarrow\;-1.4\,\mathrm{pp}$ $\uparrow\;+2.7\,\mathrm{pp}$ $\downarrow\;-2.4\,\mathrm{pp}$
Informality Rate (After) $65.5\%$ $60.8\%$ $58.5\%$ $56.3\%$
Informality Rate - $\Delta$ After-Before $\downarrow\;-2.1\,\mathrm{pp}$ $\downarrow\;-3.5\,\mathrm{pp}$ $\uparrow\;+0.4\,\mathrm{pp}$ $\downarrow\;-2.2\,\mathrm{pp}$
Note: Author’s calculations.
The correlation matrix (Figure 5) illustrates the anticipated relationships among labor market indicators. The unemployment rate and employment rate are strongly negatively correlated (-0.63), while the employment rate and participation rate exhibit a high positive correlation (0.76). As expected, the inactivity rate and participation rate demonstrate a near-perfect negative correlation (-1.0) due to their complementary definitions. With respect to informality, positive correlation with unemployment (0.50) suggests that higher joblessness is associated with increased informal employment. Conversely, the negative association with inactivity (-0.54) may indicate lower informal participation in more inactive populations. The near-zero correlation with the employment rate (0.05) suggests that the overall employment level is not directly indicative of job formality.
Granger causality tests (Figure 6), utilizing pair-specific optimal lags selected based on the Bayesian Information Criterion (BIC) ($max\leavevmode\nobreak\ lag=6$), indicate that unemployment and inactivity Granger-cause the informality rate ($p<0.05$), whereas the employment rate does not Granger-cause it ($p=0.430$). Conversely, the effect of informality on unemployment is not significant ($p=0.190$). These results are consistent across pairs, suggesting that informality primarily responds to broader labor market conditions in the short run rather than exerting a causal influence on them.
Figure 5: National labor market rates Spearman Correlation (1993-2025).
Note: Author’s calculations. Pairwise correlations between indicators were computed using Spearman’s rank correlation coefficient, which is robust to non-linear monotonic relationships and less sensitive to outliers.
Figure 6: National labor market rates Granger Causality (1993-2025).
Note: Author’s calculations. Granger causality tests were conducted allowing up to six lags, with the optimal lag for each variable pair selected via the Bayesian Information Criterion (BIC).
4.3 Departmental Labor Patterns (1993-2025)
4.3.1 Historical Analysis of Departmental Labor Market Indicators
This subsection presents the reconstructed annual trajectories of four key labor market indicators across Colombia’s 33 departments: the unemployment rate, the labor force participation rate, the employment rate, and the inactivity rate. These indicators were estimated for the period 1993-2025 using a unified methodological framework that integrates supervised learning, temporal disaggregation, and demographic harmonization. The resulting panel adheres to labor accounting identities and facilitates detailed diagnostics of both cyclical dynamics and long-term structural asymmetries across regions.
Figure 7 illustrates a diverse yet bounded distribution of unemployment rates over time. While national crises, such as the 1999 recession and the 2020 pandemic, generate noticeable transitory spikes in many departments, these events do not induce persistent unemployment divergence across the territory. Most departments exhibit unemployment rates fluctuating within a 7–14% range, with no evident outliers exceeding 20% in recent years. Departments previously highlighted, such as Guainía, Arauca, and Chocó, no longer systematically stand out; instead, transient peaks appear in various locations and dissipate over time. This observation suggests that the reconstructed series smooths local volatility while preserving the spatial-temporal imprint of systemic shocks.
Figure 7: National labor market rates (1993-2025): Yearly Unemployment (averaged) by Department
Note: Author’s calculations.
The participation rate (Figure 8) displays moderate heterogeneity across departments but converges toward a central range between 55% and 70% throughout the examined period. Contrary to prior assumptions of extreme disengagement in Amazonian or frontier regions, the updated estimates reveal more homogeneous dynamics. Departments such as Cesar, Chocó, and Caquetá exhibit gradual long-term increases in participation, with noticeable boosts following 2020. These shifts may indicate improved demographic integration, post-pandemic recovery in institutional reporting, or structural transitions in the incorporation of informal labor. Furthermore, no department consistently exceeds 75% or falls below 50%, suggesting relatively contained behavior across regions.
Figure 8: National labor market rates (1993-2025): Yearly Participation (averaged) by Department
Note: Author’s calculations.
As expected, the employment rate (Figure 9) largely aligns with the participation rate, given that unemployment remains stable across regions. The estimates confirm consistent labor absorption dynamics, with rates ranging from 45% to 65% in most departments. No regions exhibit structural employment collapses; instead, employment variation follows cyclical and symmetric patterns over time and space. Economically strong departments, such as Antioquia and Santander, display moderate rates, consistent with sectoral structures that are more stable but less intensive in informal or precarious labor.
Figure 9: National labor market rates (1993-2025): Yearly Employment (averaged) by Department
Note: Author’s calculations.
Figure 10 completes the labor identity through the lens of inactivity. The pattern aligns with participation and employment trajectories, with rates typically ranging between 30% and 45%. No department exhibits chronic detachment from the labor market. In contrast to prior interpretations that emphasized structural exclusion in regions such as Huila, Risaralda, or Sucre, the updated estimates suggest progressive reductions in inactivity and a more evenly distributed rebound following the COVID-19 shock.
Figure 10: National labor market rates (1993-2025): Yearly Inactivity (averaged) by Department
Note: Author’s calculations.
Figure 11 provides a long-term view of informality across Colombian departments. The heatmap highlights marked and persistent regional disparities. Departments such as Bogotá D.C., Valle del Cauca, Santander, and Cundinamarca consistently exhibit lower levels of informality, generally between 50% and 65%. In contrast, territories including Chocó, Vichada, La Guajira, and several Amazonian departments exhibit structurally higher informality rates, remaining above 70% throughout the period. This reflects persistent territorial gaps rather than uniform convergence. The COVID-19 shock in 2020 is evident across most departments, followed by partial recoveries at varying speeds. Some regions recovered quickly to pre-pandemic levels, while others continued to exhibit elevated informality rates, underscoring the asymmetric nature of labor market adjustments.
Figure 11: National labor market rates (1993-2025): Yearly Informality (averaged) by Department
Note: Author’s calculations.
In summary, the departmental trajectories from 1993 to 2025 reveal a labor market that, while sensitive to macroeconomic shocks, maintains a high degree of structural cohesion across regions. Unemployment fluctuations are transitory and spatially diffuse; participation and employment rates converge toward relatively narrow ranges, and inactivity shows a gradual decline with only temporary reversals during crises. Informality, although still heterogeneous, follows a long-term downward trend, exhibiting convergence across most departments with a contained rebound after COVID-19. These patterns suggest that the reconstructed series capture both the cyclical synchronization and the slow-moving structural transformations of Colombia’s labor market, providing a robust empirical foundation for diagnosing territorial asymmetries, evaluating policy impacts, and anticipating differential regional responses to future shocks.
4.3.2 Labor Market Profiles with a Quality Index
The trajectories of the EQI reveal persistent heterogeneity across Colombian departments (Figure 12). Departments such as Bogotá D.C., Santander, and Antioquia consistently achieve high scores, while Caquetá, Sucre, and Vaupés remain at the lower end of the distribution. This stratification reflects structural disparities that are not easily offset by short-term fluctuations. The penalty mechanism built into the EQI further emphasizes cases where high employment rates coexist with high informality, thereby preventing inflated assessments of fragile labor structures. Over time, the evidence indicates limited convergence: the gap between the best- and worst-performing departments has tended to persist or widen.
Relative rankings across three subperiods (1993-2003, 2004-2014, and 2015-2025) confirm these patterns (Figure 13). Bogotá D.C. and Santander retain their positions as leaders throughout, with Cundinamarca consolidating its place in the top group more recently. Conversely, Chocó, Caquetá, Arauca, and Sucre consistently occupy the lowest ranks. Some mobility is observed: Nariño and Risaralda made improvements in the most recent decade, whereas Norte de Santander and Huila declined. These movements suggest that structural disadvantages are resilient but not immutable, indicating that enhancements in labor market institutions and conditions can alter trajectories over time.
Table 8 further illustrates these disparities. The Muy bajo group, composed of seven departments, exhibits the weakest labor market conditions, characterized by persistent informality and fragile employment structures. The Bajo group, with six departments, encounters significant but less acute deficits in formality and stability. The Medio group reflects intermediate outcomes, with some departments demonstrating potential for upward mobility. The Alto group, also consisting of six departments, achieves comparatively favorable results, marked by stronger participation rates and lower informality levels. Finally, the Muy alto group, encompassing seven departments, including the largest urban economies, represents the core of high-quality employment in Colombia.
Overall, the evidence reveals a polarized pattern: a set of departments consistently positioned at the top, a group persistently at the bottom, and a middle segment with limited capacity for mobility. This persistence underscores the structural and territorial nature of labor market inequalities in Colombia.
Figure 12: Employment Quality Index (EQI) Trends (1993-2025): Monthly Smoothed Values by Department
Note: Author’s calculations.
Figure 13: Relative Ranking of Departments by EQI: Bump Chart (1993-2025)
Note: Author’s calculations.
Table 8: Classification of Colombian Departments by Employment Quality Index (EQI).
dep_code dep_name cluster_name cluster_code
18 Caquetá Muy bajo 1
27 Chocó Muy bajo 1
70 Sucre Muy bajo 1
81 Arauca Muy bajo 1
86 Putumayo Muy bajo 1
94 Guainía Muy bajo 1
97 Vaupés Muy bajo 1
20 Cesar Bajo 2
44 La Guajira Bajo 2
63 Quindío Bajo 2
73 Tolima Bajo 2
91 Amazonas Bajo 2
99 Vichada Bajo 2
17 Caldas Medio 3
23 Córdoba Medio 3
47 Magdalena Medio 3
54 Norte de Santander Medio 3
85 Casanare Medio 3
88 San Andrés, Providencia y Santa Catalina Medio 3
95 Guaviare Medio 3
08 Atlántico Alto 4
13 Bolívar Alto 4
15 Boyacá Alto 4
19 Cauca Alto 4
41 Huila Alto 4
66 Risaralda Alto 4
05 Antioquia Muy alto 5
11 Bogotá D.C. Muy alto 5
25 Cundinamarca Muy alto 5
50 Meta Muy alto 5
52 Nariño Muy alto 5
68 Santander Muy alto 5
76 Valle del Cauca Muy alto 5
Note: Author’s calculations.
4.4 Policy Analysis and Implications
National level.
The monthly panel facilitates the monitoring of both structural and cyclical labor market dynamics over more than three decades, providing a consistent benchmark for assessing countercyclical responses to crises such as those in 1999, 2008, and during COVID-19. Results indicate that informality tends to rise alongside unemployment and inactivity, highlighting the importance of integrated strategies that promote job creation while incentivizing formalization. The demographic alignment of the series ensures that labor market projections remain consistent with population aging and demographic transitions, which is essential for pension sustainability, education planning, and social protection systems.
Regional level.
The Employment Quality Index (EQI) reveals persistent heterogeneity across departments. High-performing regions (e.g., Bogotá, Antioquia, Santander, Valle del Cauca, Cundinamarca) can serve as benchmarks for the diffusion of best practices, while structurally weaker territories (e.g., Chocó, Vaupés, Arauca, Sucre, Putumayo) require targeted measures. These measures may include fiscal incentives to promote formal employment, stronger enforcement of labor regulations, and support for productive sectors with greater potential to generate stable jobs. The evidence also highlights differentiated exposure to shocks: while some regions exhibit rapid recovery, others remain persistently vulnerable. This suggests the need for tailored stabilization mechanisms, such as localized unemployment insurance schemes or emergency employment programs.
Institutional level.
The methodological framework establishes the foundation for an early-warning system in labor markets. By integrating national benchmarks, survey microdata, and machine learning-based extrapolations, this approach generates consistent estimates even in contexts with limited survey coverage. Rather than replacing primary data collection, it complements it by reducing delays and ensuring the timely availability of policy-relevant information. The transparency and replicability of the pipeline also create opportunities for adoption by statistical agencies across Latin America that face similar data constraints.
In summary, the findings emphasize the need to move beyond national averages to tackle territorial disparities in labor market outcomes. The reconstructed panel provides policymakers with the ability to identify structural deficits, design region-specific interventions, and evaluate their effectiveness with a temporal and spatial resolution that was previously unavailable.
4.5 Model Limitations and Remaining Challenges
The reconstructed labor series provide a coherent and demographically aligned panel of monthly indicators for all 33 Colombian departments from 1993 to 2025. Accounting identities are enforced, estimates are calibrated to national and departmental benchmarks, and validation confirms robust predictive performance, particularly in departments with GEIH microdata. Nevertheless, several limitations remain:
  1. 1.
    Lack of uncertainty measures: Current estimates are point predictions without intervals, limiting interpretability in data-scarce regions. Future work should incorporate Bayesian or ensemble methods to quantify uncertainty.
  2. 2.
    Indicator-by-indicator training: Labor identities are enforced only ex post. Embedding them in multi-output models could improve coherence under nonlinear shocks.
  3. 3.
    Peripheral data gaps: Amazonian, insular, and rural departments remain difficult to estimate, as dynamics are largely transferred from metropolitan areas. Strengthening primary data collection is essential.
  4. 4.
    Informality estimates: Departmental informality rates rely entirely on model-based extrapolation from urban data and thus should be interpreted as reconstructions, not observed values.
  5. 5.
    Input quality and benchmarks: Estimates depend on GEIH, demographic projections, and other sources, which may contain biases or breaks. Independent subnational benchmarks are scarce, limiting external validation.
  6. 6.
    Error heterogeneity: While national errors are low, peripheral regions show higher deviations, underscoring unequal data quality across territories.
5 Conclusions
This paper develops the first harmonized and territorially exhaustive panel of monthly labor indicators for Colombia’s 33 departments, covering the period from 1993 to 2025. The reconstruction pipeline integrates temporal disaggregation, supervised learning, deep neural architectures, retro-polarization, and demographic calibration to produce consistent series of employment, unemployment, participation, inactivity, informality, and population aggregates. The result is a high-frequency dataset with national errors below 2.5% MAPE, while maintaining demographic and labor accounting identities.
Three central findings emerge. First, supervised and deep learning models, when carefully regularized and anchored to structural proxies, can extrapolate labor indicators into territories lacking direct survey coverage. This is accomplished by leveraging signals from metropolitan markets and departmental aggregates. The neural network model demonstrated particular effectiveness in capturing heterogeneous informality dynamics that conventional econometric approaches could not replicate. This underscores the potential of modern learning methods to enhance labor statistics in contexts with incomplete territorial data.
Second, enforcing demographic and labor identities ex-post ensures numerical coherence and flexibility; however, these constraints are not inherently learned by the models. While this choice enhances predictive accuracy, it diminishes structural interpretability. Future work should explore architectures that embed these identities directly, allowing them to emerge endogenously during training.
Third, although the overall reconstruction is accurate, performance varies across space and time. Peripheral or resource-dependent departments exhibit larger errors and slower recovery following crises, whereas diversified urban economies converge more rapidly to national averages. These differences underscore the necessity of enhancing primary data collection in marginalized areas and formulating differentiated territorial policies.
Beyond technical performance, the reconstructed panel offers empirical leverage for both research and policymaking. At the national level, it reflects Colombia’s labor dynamics during significant shocks, including the 1999 financial crisis, the 2008-2009 global recession, and the COVID-19 pandemic, while maintaining internal balance. At the subnational level, it reveals persistent asymmetries: high informality and low participation in peripheral regions; low informality and high participation in diversified urban economies; and intermediate groups demonstrating mixed outcomes. Recovery trajectories also vary, with formalized regions recovering more rapidly from crises.
Methodologically, the modular pipeline provides a replicable framework for countries with urban-biased labor surveys or incomplete territorial coverage. This approach can be expanded to reconstruct additional indicators (e.g., underemployment, hours worked, or formality), incorporate sectoral and occupational decompositions using CIIU-linked auxiliary data, and adopt Bayesian or multivariate architectures to enhance uncertainty quantification. Integrating labor accounting identities directly into neural architectures represents a key frontier for improving both accuracy and interpretability.
In summary, the reconstructed panel offers a novel statistical framework for analyzing regional convergence, resilience to shocks, and informality gaps in Colombia. Although it does not serve as a replacement for direct statistical production, it complements official sources by addressing territorial and temporal gaps, facilitating retrospective diagnostics, and informing policy strategies for more inclusive labor markets.
Ethical Considerations
This study relies exclusively on aggregated, anonymized, and publicly available data sourced from official statistical agencies, including the Departamento Administrativo Nacional de Estadística (DANE), the World Bank, and the International Labour Organization (ILO). At no stage were individual-level, confidential, or personally identifiable records accessed, stored, or processed. No ethical approval was required due to the non-sensitive nature of the data. The study adheres to the principles of research transparency and reproducibility by documenting all methodological steps and ensuring the open availability of derived results.
Acknowledgments
The author acknowledges the institutional efforts of DANE, ILOSTAT, and the World Bank in providing unrestricted access to labor market and demographic statistics. The views, interpretations, and conclusions expressed in this study are solely those of the author.
Funding and Competing Interests
This research did not receive funding from public agencies, private institutions, or nonprofit organizations. The author declares no competing interests, whether financial, professional, or personal.
Data Availability Statement
All reconstructed labor market indicators generated in this study, including monthly national and departmental series from 1993 to 2025, are publicly available at: https://doi.org/10.5281/zenodo.16899686
The repository contains machine-readable data files, methodological documentation, and version-controlled metadata. All input data used in the reconstruction process, including GEIH labor annexes, World Bank indicators, and ILOSTAT aggregates, are publicly accessible through their respective institutional repositories.
Code Availability Statement
All source code utilized for data processing, reconstruction, and visualization is publicly available in the GitHub repository archived on Zenodo: https://doi.org/10.5281/zenodo.16899686 (also accessible at https://github.com/jaimevera1107/colombian-labor-market)
The codebase is version-controlled, thoroughly documented, and comprises reproducible scripts necessary for replicating all analyses and figures presented in this paper.
Appendix Appendix A Institutional Context of Labor Statistics in Colombia
Appendix A.1 National Labor Information Infrastructure
Labor statistics in Colombia are produced primarily by the Departamento Administrativo Nacional de Estadística (DANE). The main source is the Gran Encuesta Integrada de Hogares (GEIH), a continuous household survey that provides monthly estimates of employment, unemployment, and labor force participation. Although nationally representative, the GEIH encounters ongoing challenges related to territorial and temporal coverage, which restrict its utility for subnational policy analysis.
Appendix A.2 Scope and Design of the GEIH
The GEIH was launched in 2006 to unify previous household surveys. Its initial design generated monthly estimates for Bogotá and 23 departmental capitals, excluding peripheral regions such as the Amazon and Orinoquía. In 2019, the sample was expanded to cover all 32 departmental capitals in addition to Bogotá. However, this expansion remains urban-focused; rural areas are surveyed quarterly and are typically aggregated into annual estimates, which results in an underrepresentation of rural labor dynamics.
Appendix A.3 Regional Gaps and Statistical Coverage
Despite Colombia’s decentralized governance, which assigns responsibilities for employment and development planning to departments, statistical coverage does not align with these mandates. Differences in population size, informality, and economic structure (e.g., Antioquia vs. Vaupés) are not adequately captured, restricting the ability to monitor regional labor markets and evaluate place-based policies.
Appendix A.4 Institutional Fragmentation
Besides DANE, other entities generate labor-related data, including the Ministry of Labor, the Banco de la República, and the Departamento Nacional de Planeación (DNP). However, integration across these sources is limited, particularly at disaggregated territorial and temporal levels. Moreover, DANE’s regular outputs emphasize urban areas, necessitating indirect mapping from city data to departmental levels. It is important to note that coverage of all departmental capitals does not ensure comprehensive representativeness across departments, particularly in dispersed regions.
Appendix Appendix B Data Preprocessing and National Baseline Reconstruction
Appendix B.1 Interpolation of Annual Series
ILO and World Bank indicators for employment, unemployment, and participation were available only on an annual basis between 1993 and 2000. These indicators were interpolated using the Akima method (cubic Hermite spline) to avoid spurious oscillations. Interpolation was conducted on rates prior to deriving levels, ensuring smooth transitions across the baseline.
Appendix B.2 Smoothing of Rates
To mitigate short-term volatility, centered moving averages were applied selectively. In particular, the unemployment rate (2007–2016) was smoothed to reduce fluctuations associated with the global financial crisis and the oil price shock. Smoothing was restricted to series exhibiting significant high-frequency noise to preserve structural dynamics.
Appendix B.3 Participation Rate Reconstruction
The participation rate ($PR_{t}$) was reconstructed using local interpolation and smoothing, adjusted with a scaling factor ($\alpha_{t}$) defined as the median labor force-to-working-age population ratio during stable periods. This approach ensured long-term demographic alignment and corrected for discontinuities in the original sources.
Appendix B.4 Conversion to Levels and Accounting Closure
Smoothed rates were converted into proportions relative to the working-age population ($PET_{t}$). Employment ($E_{t}$), unemployment ($U_{t}$), inactivity ($I_{t}$), and labor force ($PEA_{t}$) were derived from accounting identities:
$PET_{t}=E_{t}+U_{t}+I_{t},\quad PEA_{t}=E_{t}+U_{t}$
Residuals were required to remain within ±0.1%. Deviations exceeding this threshold triggered diagnostic checks and subsequent recalibration. Departmental reconstructions were then aggregated to the national level and compared with the baseline.
Appendix B.5 Error Metrics
Model accuracy was evaluated using Mean Absolute Error (MAE), Root Mean Square Error (RMSE), and Absolute Percentage Error (APE) to assess both absolute and relative deviations. These metrics were computed for both national and departmental aggregates.
Appendix B.6 Disaggregation Strategy
Annual labor ratios $y_{t}$ were disaggregated into monthly series $\hat{y}_{tm}$ using the TempDisaggModel (Chow-Lin-Opt, average conversion), with monthly indicators $x_{tm}$ (CPI, TRM, real wages, PPI) contributing to intra-annual variation:
$\hat{y}_{tm}=\text{TD}(y_{t},x_{tm};\theta)$
Annual re-aggregation verified the consistency with the original totals, and the relative RMSE against the GEIH monthly series was utilized to select the optimal indicator $x^{*}$ for each variable.
Appendix B.7 Rescaling and Adjustment
The selected monthly ratios $\rho_{tm}$ were rescaled to levels based on DANE’s annual working-age population:
$\text{Level}_{tm}=\rho_{tm}\cdot PET_{t}$
Accounting identities were maintained through residual closure:
$\text{PEA}_{tm}=\text{Employed}_{tm}+\text{Unemployed}_{tm},\quad\text{PET}_{tm}=\text{Labor Force}_{tm}+\text{Inactive}_{tm}$
PET series were linearly interpolated to prevent artificial jumps, and the residuals were graphically monitored to validate coherence.
Appendix B.8 Validation Outputs
The relative RMSE of each indicator-ratio pair is reported in Table B.1. The lowest-error proxy for each labor variable is highlighted, which confirms the robustness of the disaggregation.
Table B.1: Relative RMSE by Labor Ratio and Monthly Indicator
Ratio CPI TRM Deflated Wage PPI Best Indicator
Employment 0.0144 0.0265 0.0180 0.0150 CPI
Unemployment 0.0068 0.0074 0.0063 0.0069 Deflated Wage
PEA 0.0140 0.0287 0.0161 0.0151 CPI
PET 0.0109 0.0382 0.0165 0.0149 CPI
Inactivity 0.0104 0.0158 0.0114 0.0114 CPI
Note: Author’s calculations based on harmonized World Bank-ILO estimates and official GEIH series from DANE. RMSE values are computed by comparing annualized disaggregated series to GEIH national estimates.
The figures presented below, generated by the automated pipeline, provide graphical validations:
  • Figure B.1 illustrates the reconstructed labor ratios over time.
  • Figure B.2 depicts the resultant absolute levels following the rescaling.
  • Figure B.3 shows the residual deviations from the labor market identities after the closure adjustment.
Figure B.1: Disaggregated Monthly Labor Ratios (Pre-Closure)
Note: Author’s calculations. Monthly ratios obtained from the disaggregation process prior to rescaling and identity adjustments.
Figure B.2: Final Disaggregated Labor Series (Post-Closure, Absolute Values)
Note: Author’s calculations. Absolute monthly levels obtained after rescaling labor ratios with official population projections and applying accounting closure.
Figure B.3: Residuals from Labor Identities after PET Smoothing and Closure
Note: Author’s calculations. Residual gaps in PET and PEA identities, capturing misfits introduced during the smoothing and closure steps.
Appendix Appendix C Construction of City Monthly Labor Series
This appendix describes the methodology used to reconstruct harmonized monthly labor indicators at the city level. The goal is to obtain a panel $y_{c,t,m}$ that is temporally consistent, statistically coherent, and aligned with both departmental and national benchmarks. Cities serve as the primary high-frequency units of observation, capturing intra-annual dynamics that are subsequently scaled to ensure compatibility with higher levels of aggregation.
Appendix C.1 Coverage and Notation
Let $C$ denote the set of cities observed in GEIH, with each city $c\in C$ belonging to one department $d\in D$. Indicators of interest are grouped in the set $\mathcal{V}=\{\text{employed},\text{inactive},\text{pea},\text{pet},\text{population},\text{unemployed}\}$. Each observation is indexed by year $t$ and month $m\in\{1,\dots,12\}$.
Appendix C.2 Similarity Mapping and Urban Proxies
For cities directly covered by GEIH, monthly indicators $v^{\text{obs}}_{c,t,m}$ are available. For cities or departments with incomplete coverage, similarity mapping is employed. Correlation coefficients across overlapping periods are computed between partially observed series and candidate reference cities. For each target $c$, the closest counterpart is selected:
$c^{*}(v)\;=\;\arg\max_{j\in C}\rho^{(c,j)}_{v}$,
where $\rho^{(c,j)}_{v}$ denotes the rank correlation of indicator $v$. This ensures that missing series inherit the dynamics of the most similar urban counterpart.
Appendix C.3 Synthetic Annual Alignment
Once urban signals are obtained, they are aggregated at the annual level to maintain coherence with city-year benchmarks:
$\tilde{v}_{c,t}\;=\;\frac{1}{12}\sum_{m=1}^{12}v^{\text{obs}}_{c,t,m}.$
When annual reference values $v_{c,t}$ are available, proportional splicing is employed:
$\hat{v}_{c,t,m}\;=\;v^{\text{obs}}_{c,t,m}\cdot\frac{v_{c,t}}{\tfrac{1}{12}\sum_{m=1}^{12}v^{\text{obs}}_{c,t,m}}\quad\text{for ratios,}$
with the denominator replaced by the annual sum for level variables. This ensures that monthly city profiles either sum or average to the known annual totals.
Appendix C.4 Temporal Disaggregation of Annual Signals
For cases where only annual city totals $v_{c,t}$ are observed, Chow-Lin temporal disaggregation with AR(1) errors is employed:
$\mathbf{y}_{c}=\mathbf{C}\mathbf{x}_{c}+\boldsymbol{\varepsilon}_{c},\qquad\text{Cov}(\varepsilon_{t},\varepsilon_{t^{\prime}})=\sigma^{2}\rho^{|t-t^{\prime}|},$
where $\mathbf{y}_{c}$ denotes annual targets, $\mathbf{x}_{c}$ represents monthly national indicators aggregated through the conversion matrix $\mathbf{C}$, and $\rho$ captures serial correlation. This method allocates annual city totals across months in a manner that aligns with both national dynamics and city-specific profiles.
Appendix C.5 National Alignment
After disaggregation, proportional adjustment ensures that the total across all cities aligns with the national monthly benchmarks:
$\bar{v}_{c,t,m}=v^{\text{nat}}_{t,m}\cdot\frac{\hat{v}_{c,t,m}}{\sum_{c^{\prime}\in C}\hat{v}_{c^{\prime},t,m}},$
where $v^{\text{nat}}_{t,m}$ are national monthly estimates. This ensures additivity among the city, departmental, and national layers.
Appendix C.6 Accounting Closure and Demographic Smoothing
To maintain internal coherence, labor accounting identities are established:
$\text{unemployed}_{c,t,m}=\text{pea}_{c,t,m}-\text{employed}_{c,t,m},\qquad\text{pet}_{c,t,m}=\text{pea}_{c,t,m}+\text{inactive}_{c,t,m}.$
Population data is interpolated to ensure smooth trajectories, and PET is linearly adjusted between annual anchors to prevent artificial discontinuities. Smoothing procedures ensure that demographic series evolve gradually over time while remaining consistent with annual benchmarks.
Appendix Appendix D Reconstruction of Department Monthly Labor Series
This appendix describes two complementary procedures for reconstructing monthly labor indicators at the departmental level. Let $D$ denote the set of all departments. Let $D_{\text{city}}\subset D$ represent the departments with monthly city-level observations from GEIH annexes, while $D_{\text{missing}}=D\setminus D_{\text{city}}$ signifies the remaining departments. The objective is to obtain monthly ratios and levels that are internally consistent, temporally coherent, and aligned with national benchmarks.
Appendix D.1 Reconstruction Based on Monthly City-Level Observations
Similarity mapping.
For each target department $d\in D_{\text{missing}}$ and each labor ratio $y\in\mathcal{Y}$, we compute Spearman correlations between the monthly GEIH series of $d$’s candidate counterparts in $D_{\text{city}}$ and the available annual or partial monthly data of $d$, within the largest common time window. The most similar department is
$d^{*}(y)\;=\;\arg\max_{j\in D_{\text{city}}}\rho^{(d,j)}_{y}$,
where $\rho^{(d,j)}_{y}$ represents the Spearman correlation for the ratio $y$.
Synthetic monthly path.
The initial synthetic monthly path for $d$ is derived from its optimal counterpart:
$\tilde{y}_{d,t,m}\;=\;y^{\text{GEIH}}_{d^{*}(y),\,t,m},$
with $t$ year and $m\in\{1,\dots,12\}$. This approach maintains the intra-annual dynamics observed in the GEIH cities.
Annual splicing to preserve departmental aggregates.
To align with established annual departmental benchmarks $y_{d,t}$, we implement proportional splicing each year:
$\hat{y}_{d,t,m}\;=\;\tilde{y}_{d,t,m}\cdot\frac{y_{d,t}}{\frac{1}{12}\sum_{m=1}^{12}\tilde{y}_{d,t,m}}\quad\text{for ratios (average rule).}$
For levels, replace the denominator by $\sum_{m=1}^{12}\tilde{y}_{d,t,m}$ (sum rule).
National alignment (monthly proportional adjustment).
Given national monthly benchmarks $y^{\text{nat}}_{t,m}$, we ensure consistency in totals by applying proportional adjustments:
$\bar{y}_{d,t,m}\;=\;y^{\text{nat}}_{t,m}\cdot\frac{\hat{y}_{d,t,m}}{\sum_{d^{\prime}\in D}\hat{y}_{d^{\prime},t,m}}.$
Accounting closure and demographic smoothing.
Using monthly ratios and the working-age population, we derive levels and enforce identities:
$\text{Unemployed}_{d,t,m}$ $=\text{PEA}_{d,t,m}-\text{Employed}_{d,t,m},$
$\text{PET}_{d,t,m}$ $=\text{PEA}_{d,t,m}+\text{Inactive}_{d,t,m}.$
Monthly PET and population trajectories are linearly interpolated between yearly anchors to eliminate spurious jumps. Residuals from $\text{PEA}=\text{Employed}+\text{Unemployed}$ and $\text{PET}=\text{PEA}+\text{Inactive}$ are monitored and maintained within small tolerances.
Output.
We obtain monthly indicators $y_{d,t,m}$ for all $y\in\mathcal{Y}$ and $d\in D_{\text{city}}\cup D_{\text{missing}}$, which are consistent with departmental annual benchmarks and national monthly totals.
Appendix D.2 Temporal Disaggregation of Annual Departmental Indicators
Objective and scope.
For departments that utilize only annual indicators $y_{d,t}$, we estimate monthly series $\hat{y}_{d,t,m}$ such that
$\frac{1}{12}\sum_{m=1}^{12}\hat{y}_{d,t,m}\;=\;y_{d,t}\quad\text{for ratios (average rule),}$
and $\sum_{m=1}^{12}\hat{y}_{d,t,m}=y_{d,t}$ for levels (sum rule).
Model specification (Chow-Lin with AR(1) errors).
We employ the Chow-Lin method, utilizing average conversion and an autoregressive (AR(1)) error structure:
$\mathbf{y}_{d}\;=\;\mathbf{C}\mathbf{x}_{d}+\boldsymbol{\varepsilon}_{d},\qquad\text{Cov}(\varepsilon_{d,t},\varepsilon_{d,t^{\prime}})=\sigma^{2}\rho^{|t-t^{\prime}|},$
where $\mathbf{y}_{d}$ represents annual targets, $\mathbf{x}_{d}$ denotes monthly indicators aggregated by the average rule via $\mathbf{C}$, and $\rho$ is estimated within the disaggregation routine.
Input data format.
We organize inputs in a long format that is compatible with the disaggregation pipeline:
  • Index: calendar year $t$ (integer).
  • Grain: month $m\in\{1,\dots,12\}$ (integer).
  • y: annual departmental target for variable $v$.
  • X: selected national monthly indicator for $v$ (e.g., CPI, TRM, deflated wage, PPI).
  • dep_code: department identifier.
National alignment and closure.
After disaggregation, we apply the same monthly proportional adjustment to match national totals and enforce labor accounting identities as described above, using linear interpolation on PET to ensure demographic coherence.
Appendix Appendix E Construction of Economic Indicators
To establish a consistent macroeconomic and institutional framework for the supervised estimation of departmental labor indicators, we constructed a harmonized monthly dataset of national-level variables. These variables serve as macro-institutional control signals, ensuring demographic and economic coherence across departmental projections. The inclusion criteria were: (i) availability at a monthly frequency; (ii) consistent publication since 1993; and (iii) direct relevance to labor market dynamics.
Appendix E.1 Macroeconomic Variables
The baseline macroeconomic indicators were sourced from Banco de la República and DANE. These indicators include:
  • TRM (TRM, Tasa Representativa del Mercado), Monthly average of the representative exchange rate between the Colombian peso and the US dollar.
  • CPI (IPC, Índice de Precios al Consumidor), Headline monthly inflation index and its 12-month variation.
  • PPI (IPP, Índice de Precios al Productor), Producer price index, capturing upstream cost dynamics.
  • X (FOB, Exportaciones), Monthly value of total Colombian exports (in USD).
  • M (CIF, Importaciones), Monthly value of total Colombian imports (in USD).
  • UVR (UVR, Unidad de Valor Real), Monthly value of the UVR, indexed to inflation expectations and relevant for financial contracts.
  • IPI (IPI, Índice de Producción Industrial), Monthly industrial production index, published by DANE as a proxy for real sector activity.
All series were processed to ensure temporal continuity. Missing months were reconstructed primarily using spline-based methods (Akima/PCHIP). Simple forward fill was restricted to gaps of $\leq$12 months to avoid bias and step artifacts. Consistency was validated against official annual summaries.
Appendix E.2 Institutional Wage Components
To capture labor market institutions that directly affect income floors, we incorporated two key statutory components:
  • Minimum Legal Monthly Wage (SMMLV, Salario Mínimo Mensual Legal Vigente), Annual Nominal statutory minimum wage in force each January, extended uniformly across months.
  • Transport Subsidy (ST, Subsidio de Transporte), Annual subsidy value granted to workers earning up to two minimum wages, updated annually.
Both were obtained from official decrees and labor bulletins. To express their joint real value, we constructed a deflated composite salary index, using 2008 as the base year ($2008=100$):
$\text{Deflated Salary}_{t}=\frac{\text{Minimum Wage}_{t}+\text{Transport Subsidy}_{t}}{\text{CPI}_{t}}\times 100$
This index approximates the purchasing power of statutory income floors, adjusted for inflation. It was preferred over separate indicators as it better reflects the effective statutory floor encountered by low-income workers.
Appendix E.3 Clusters of Departments
To apply structural smoothing in departmental estimates, we grouped departments into clusters defined by the size of the urban population, the degree of labor market consolidation, and the historical production structure. These clusters do not replace individual values; instead, they guide smoothing in cases where series exhibit high volatility or limited coverage.
Table E.2: Department Clusters Used for Structural Smoothing
Department Code Assigned Cluster
05 (Antioquia) large_urban
08 (Atlántico) large_urban
11 (Bogotá D.C.) large_urban
25 (Cundinamarca) large_urban
76 (Valle del Cauca) large_urban
68 (Santander) consolidated
73 (Tolima) consolidated
50 (Meta) consolidated
70 (Sucre) consolidated
44 (La Guajira) peripheral
27 (Chocó) peripheral
81 (Arauca) peripheral
91 (Amazonas) peripheral
94 (Guainía) peripheral
95 (Guaviare) peripheral
97 (Vaupés) peripheral
99 (Vichada) peripheral
52 (Nariño) atypical_strong
85 (Casanare) atypical_strong
13 (Bolívar) mid_urban
20 (Cesar) mid_urban
23 (Córdoba) mid_urban
41 (Huila) mid_urban
47 (Magdalena) mid_urban
15 (Boyacá) mid_urban
18 (Caquetá) mid_urban
19 (Cauca) mid_urban
54 (Norte de Santander) mid_urban
63 (Quindío) mid_urban
66 (Risaralda) mid_urban
17 (Caldas) mid_urban
Note: Author’s own construction.
Appendix Appendix F Data Integration
Appendix F.1 Temporal Alignment and Integration
All economic and institutional variables were compiled into a single monthly panel covering the period from January 1993 to June 2025. Each observation includes a year and month field to facilitate integration with the modeling dataset.
Where necessary, all monetary series were converted to real terms using the Consumer Price Index (CPI) with a base of 100 established in January 1993. The resulting panel was standardized and merged with the prediction dataset by calendar time, ensuring consistent availability of covariates for each departmental-year-month combination modeled.
This harmonized panel establishes the economic foundation for the predictive framework and is utilized during both the training and projection phases.
Appendix F.2 Data preparation for estimation
For departments with direct GEIH observations, we reconstructed complete monthly series back to 1993 using a proportional backward splicing method. Let $t^{*}$ denote the earliest month with valid GEIH data for department $d$ and variable $k$. When a sufficiently strong correlation was observed between $y_{d,t}^{(k)}$ (departmental series) and a city-level or cluster proxy $z_{d,t}^{(k)}$, missing values for $t<t^{*}$ were estimated as follows:
$\hat{y}_{d,t}^{(k)}=y_{d,t^{*}}^{(k)}\cdot\frac{z_{d,t}^{(k)}}{z_{d,t^{*}}^{(k)}},\quad t<t^{*}$
This assumes a stable proportional relationship between departmental and proxy dynamics over time. In instances where this correlation was weak, no splicing was applied, and the series was left unextended.
Labor accounting identities were established to ensure internal consistency and adjust for marginal imbalances:
$\text{PEA}_{d,t}=\text{Employed}_{d,t}+\text{Unemployed}_{d,t}$
$\text{Inactives}_{d,t}=\text{PET}_{d,t}-\text{PEA}_{d,t}$
$\text{PET}_{d,t}\leq\text{Population}_{d,t}$
Appendix Appendix G Supervised Estimation Methodology
Appendix G.1 Estimation Procedure
Let $d\in\{1,\dots,D\}$ denote departments, and $t\in\{1,\dots,T\}$ denote monthly time periods from January 1993 to June 2025. For each $(d,t)$ pair, we aim to estimate absolute labor indicators $\{y_{d,t}^{(j)}\}_{j=1}^{J}$, where $j$ indexes five labor categories: PET, PEA, employed, unemployed, and inactive population.
Prior to modeling, all targets $y_{d,t}^{(j)}$ are scaled by their respective population divisors $N_{d,t}$ to obtain relative shares:
$r_{d,t}^{(j)}=\frac{y_{d,t}^{(j)}}{N_{d,t}}.$
This step ensures comparability across departments of different sizes and allows conversion to absolute values following estimation.
Let $\mathbf{x}_{d,t}\in\mathbb{R}^{p}$ be the feature vector at time $t$ for department $d$, which includes:
  • Population features (e.g., $N_{d,t}$, urban/rural composition).
  • National monthly macroeconomic indicators.
  • City-level labor indicators from GEIH, available only in observed departments.
The estimator $f_{\theta}:\mathbb{R}^{p}\to\mathbb{R}^{J}$ is designed to predict all $J$ target rates concurrently:
$\hat{\mathbf{r}}_{d,t}=f_{\theta}(\mathbf{x}_{d,t}).$
We employ a multi-output framework, utilizing a distinct XGBoost model for each target variable. Each model is trained on the observed departments, incorporating standardized features and target shares. Input and output scalers are fitted solely on the training data to prevent information leakage. No hyperparameter tuning was conducted; default XGBoost parameters were consistently utilized across all targets.
Appendix G.2 Evaluation of Model Performance
The predictive performance of each model is evaluated using four complementary validation schemes:
  1. 1.
    Full-sample fit (no holdout): The model is trained on the complete dataset without any validation split. This provides an optimistic reference for in-sample predictive accuracy and is used as the final production model for extrapolation.
  2. 2.
    Holdout validation: An 80-20 split stratified by year is applied to simulate out-of-sample forecasting. The model is trained on 80% of the data and tested on the remaining 20%, ensuring temporal representativeness in both sets.
  3. 3.
    Leave-K-Out Cross-Validation (LKO): A robust generalization test is performed by randomly excluding $k=5$ departments in each of $n=10$ iterations. This assesses the model’s capacity to generalize to regions partially excluded from the training set. For each iteration, training is conducted on the remaining departments, and metrics are aggregated.
  4. 4.
    Leave-One-Group-Out (LOGO): In this strictest validation setup, each department is excluded once and used as the test set. This directly evaluates extrapolation capability to entirely unobserved spatial units.
In all validation schemes, population-based scaling is applied to both inputs and targets prior to model fitting. After prediction, all outputs are reverted to absolute levels by multiplying the predicted rates by departmental population estimates.
$\hat{y}_{d,t}=\hat{r}_{d,t}\cdot N_{d,t}.$
For each validation fold and each labor indicator, we calculate the following error metrics:
  • Root Mean Squared Error (RMSE): quantifies the absolute prediction error in individual units.
  • Mean Absolute Error (MAE): provides a robust measure of average deviation.
  • Mean Absolute Percentage Error (MAPE): contextualizes the error as a percentage of the true value.
Let $y_{d,t}$ denote the observed value and $\hat{y}_{d,t}$ denote the predicted value. Then:
$\text{MAPE}_{d,t}=\left|\frac{y_{d,t}-\hat{y}_{d,t}}{y_{d,t}}\right|\times 100$
All metrics are aggregated annually and by target indicator to evaluate temporal consistency and differential accuracy across labor variables. The final results for each validation strategy are presented in Table G.3.
Appendix G.3 Post-Processing and Calibration
Predicted levels are adjusted to align with the following criteria:
$\text{Employed}_{d,t}+\text{Unemployed}_{d,t}$ $=\text{PEA}_{d,t},$
$\text{PEA}_{d,t}+\text{Inactive}_{d,t}$ $=\text{PET}_{d,t},$
$\text{PET}_{d,t}$ $\leq\text{Population}_{d,t}.$
Residuals are allocated proportionally when necessary. Subsequently, proportional scaling aligns departmental aggregates with national benchmarks:
$\hat{y}_{d,t}^{\text{cal}}=\hat{y}_{d,t}\cdot\frac{\hat{Y}_{t}^{\text{nat}}}{\sum_{d}\hat{y}_{d,t}}.$
This ensures coherence between departmental estimates and previously reconstructed national totals.
Appendix G.4 Supplementary materials
Table G.3: Validation Performance Summary Across All Labor Indicators (share scale)
Validation Scheme MAE RMSE MAPE (%)
Holdout (stratified) 0.0029 0.0048 1.3761
Holdout (random 80-20) 0.0081 0.0114 3.7621
Leave-K-Out (K=3, N=20) 0.0303 0.0417 10.2179
Leave-One-Out (unique) 0.0301 0.0383 10.6621
Note: Author’s calculations. The Mean Absolute Percentage Error (MAPE) is expressed as the mean absolute percentage error across all department-month-target combinations.
Figure G.4: Distribution of MAPE in model performance for LOO strategy
Note: Author’s calculations. Distribution of metrics across Leave-One-Out.
Appendix G.5 Comparison of Main Labor Market Results
Figure G.5: Annual observed vs. estimated values of employed population by department
Note: Author’s calculations. Solid lines represent official GEIH aggregates; dashed lines are model estimates.
Figure G.6: Annual observed vs. estimated values of unemployed population by department
Note: Author’s calculations. Solid lines represent official GEIH aggregates; dashed lines are model estimates.
Figure G.7: Annual observed vs. estimated values of inactive population by department
Note: Author’s calculations. Solid lines represent official GEIH aggregates; dashed lines are model estimates.
Figure G.8: Annual observed vs. estimated values of economically active population (PEA) by department
Note: Author’s calculations. Solid lines represent official GEIH aggregates; dashed lines are model estimates.
Figure G.9: Annual observed vs. estimated values of working-age population (PET) by department
Note: Author’s calculations. Solid lines represent official GEIH aggregates; dashed lines are model estimates.
Figure G.10: Annual observed vs. estimated values of total population by department
Note: Author’s calculations. Solid lines represent official GEIH aggregates; dashed lines are model estimates.
Appendix Appendix H Informality Estimation with Neural Networks
Appendix H.1 Formal Specification of the Neural Model and National Reconciliation
Appendix H.1.1 Notation and Preprocessing
Let
$\mathcal{D}=\{(\mathbf{x}_{t,g},y_{t,g},d_{g},\tau_{t})\}_{t=1,\dots,T;\ g\in\mathcal{G}},$
where $\mathbf{x}_{t,g}\in\mathbb{R}^{p}$ denotes the covariates, $y_{t,g}\in(0,1)$ the informality rate, $d_{g}$ the department code, and $\tau_{t}$ a monthly index.
Before training, covariates are transformed column-wise using a configurable scaler $S(\cdot)$, which may include standardization, min-max normalization, or robust scaling. This transformation is consistently applied to the training set to prevent data leakage.
Target transform.
The logit of the informality rate is modeled to constrain predictions within $(0,1)$:
$z=\phi(y)=\log\!\left(\frac{y}{1-y}\right),\qquad y=\phi^{-1}(z)=\sigma(z)=\frac{1}{1+e^{-z}}.$
Appendix H.1.2 Residual MLP Backbone
The network is a residual multilayer perceptron that features configurable hidden dimensions, number of blocks, and an adjustable dropout rate. Each block encompasses Layer Normalization, a linear transformation, dropout, and ReLU activation, while incorporating residual connections when enabled.
In the baseline configuration, the hidden dimension is set to 64, dropout is set to 0.2, and two residual blocks are incorporated. The final output layer produces one scalar per observation, interpreted as the logit of the predicted informality rate.
Appendix H.1.3 Optimization
Training minimizes the Huber loss in logit space with parameter $\delta=1$. The optimizer employed is Adam, with a learning rate of $\eta=10^{-3}$, weight decay of $\lambda_{\mathrm{wd}}=10^{-4}$, and an adaptive batch size capped at $512$. Early stopping is implemented with a patience of =20=20 and a minimum improvement of $\Delta_{\min}=10^{-4}$. Validation splits are performed temporally when a date column is available. A fixed random seed is utilized to ensure reproducibility.
Appendix H.1.4 Departmental Reconciliation to National Total
Let $\hat{I}_{d,t}$ denote the departmental rate, and $E_{d,t}$ represent the departmental employment. Compute the implied informal counts $\hat{U}_{d,t}=\hat{I}_{d,t}\,E_{d,t}$ and reconcile them with the national benchmark through proportional scaling:
$\lambda_{t}=\frac{I_{t}^{\text{nat}}\sum_{d}E_{d,t}}{\sum_{d}\hat{U}_{d,t}},\qquad\hat{U}_{d,t}^{\text{cal}}=\lambda_{t}\,\hat{U}_{d,t},$
$\hat{I}_{d,t}^{\text{cal}}=\frac{\hat{U}_{d,t}^{\text{cal}}}{E_{d,t}}.$
This ensures $\sum_{d}\hat{U}_{d,t}^{\text{cal}}=I_{t}^{\text{nat}}\sum_{d}E_{d,t}$ for each $t$, while maintaining cross-departmental variation.
Appendix H.1.5 Validation Protocols
Let $\mathcal{E}\in\{\mathrm{RMSE},\mathrm{MAE},\mathrm{MAPE}\}$:
$\mathrm{RMSE}=\sqrt{\tfrac{1}{n}\sum_{i}(y_{i}-\hat{y}_{i})^{2}},\quad\mathrm{MAE}=\tfrac{1}{n}\sum_{i}|y_{i}-\hat{y}_{i}|,\quad\mathrm{MAPE}=100\cdot\tfrac{1}{n}\sum_{i}\frac{|y_{i}-\hat{y}_{i}|}{\max(|y_{i}|,\varepsilon)}.$
We evaluate using post-cutoff data under the following conditions:
Table H.4: Performance metrics for the $N$ model under different validation protocols (post-2007 period).
Validation scheme RMSE MAE MAPE (%)
Full 0.0149 0.0111 2.0045
80/20 0.0169 0.0123 2.1938
LOO 0.0716 0.0543 9.8615
LKO 0.0723 0.0556 10.6612
Note: RMSE and MAE are expressed in proportion units, MAPE in percentage terms. Evaluations restricted to post-2007 data to ensure comparability.
Figure H.11: Observed vs. estimated monthly informality rates at the city level.
Note: Author’s calculations. Solid lines represent official GEIH aggregates; dashed lines are model estimates.
Appendix Appendix I Employment Quality Index (EQI)
Let $d\in\mathcal{D}$ denote departments and $t\in\mathcal{T}$ months. For each $(d,t)$ observation, the Employment Quality Index (EQI) is constructed through the following steps:
  • Indicators. The EQI integrates four core labor market indicators:
    • Employment rate ($E$)
    • Unemployment rate ($U$)
    • Inactivity rate ($I$) / Participation rate ($P$)
    • Informality rate ($N$) / Formality rate ($F$)
  • Smoothing. Each indicator series $m_{d,t}$ is smoothed using a centered moving average with window length $w=13$:
    $\tilde{m}_{d,t}=\frac{1}{|\mathcal{W}|}\sum_{\tau\in\mathcal{W}(t)}m_{d,\tau},\quad\mathcal{W}(t)=\{t-6,\ldots,t+6\}.$
    Missing edges are filled by carrying forward the nearest available values.
  • Winsorization. To reduce the influence of outliers, the smoothed series $\tilde{m}_{d,t}$ is truncated at the 1st and 99th percentiles:
    $\tilde{m}_{d,t}^{*}=\min\{\max\{\tilde{m}_{d,t},q_{\alpha}\},q_{1-\alpha}\},\quad\alpha=0.01.$
  • Normalization. Indicators are transformed into percentile ranks. By default, this is done within each month $t$ (preferred, to avoid level shifts), but can also be applied globally across all $t\in\mathcal{T}$. For indicator $m$:
    $r_{d,t}^{(m)}=\frac{\text{rank}(\tilde{m}_{d,t}^{*})}{|\mathcal{D}|}.$
    Positively oriented indicators (e.g., $E,P,F$) retain their rank:
    $s_{d,t}^{(m)}=r_{d,t}^{(m)},$
    while negatively oriented indicators (e.g., $U,I,N$) are inverted:
    $s_{d,t}^{(m)}=1-r_{d,t}^{(m)}.$
  • Composite Index. With normalized weights $w_{m}$ satisfying $\sum_{m\in\mathcal{M}}w_{m}=1$, the baseline EQI is defined as:
    $\text{EQI}_{d,t}^{0}=100\times\sum_{m\in\mathcal{M}}w_{m}\,s_{d,t}^{(m)}.$
  • Conflict Penalty. To avoid overestimation when favorable and unfavorable conditions coexist (e.g., high $E$ with high $N$), a penalty factor $\pi_{d,t}\in[0,1]$ is applied. For a conflict pair $(a,b)$ with raw percentiles $(r_{d,t}^{(a)},r_{d,t}^{(b)})$ and penalty strength $\lambda\in[0,1]$:
    $\pi_{d,t}=1-\lambda\,r_{d,t}^{(a)}r_{d,t}^{(b)}.$
    The penalized EQI is:
    $\text{EQI}_{d,t}=\pi_{d,t}\cdot\text{EQI}_{d,t}^{0}.$
  • Clustering. Departments are classified into $K$ ordered groups based on the empirical quantiles of their long-run average:
    $\overline{\text{EQI}}_{d}=\frac{1}{|\mathcal{T}|}\sum_{t\in\mathcal{T}}\text{EQI}_{d,t}.$
    The groups $\{G_{1},\ldots,G_{K}\}$ are labeled from lowest to highest quality (e.g., Muy Bajo, Bajo, Medio, Alto, Muy Alto).
  • Ranking and Visualization. For custom periods $\mathcal{T}_{p}=[y_{0},y_{1}]$, average EQI values are computed as:
    $\text{EQI}_{d,p}=\frac{1}{|\mathcal{T}_{p}|}\sum_{t\in\mathcal{T}_{p}}\text{EQI}_{d,t}.$
    Departments are ranked within each period $p$, and their trajectories are visualized using bump charts, highlighting persistence, mobility, and divergence across time.
Appendix Appendix J Data Sources
Appendix J.1 External National labor market indicators
Table J.5: International Labor Indicators Used for National Benchmark Reconstruction
Indicator Source Frequency Coverage
Employed Population World Bank - WDI Annual 1991-2024
Unemployed Population World Bank - WDI Annual 1991-2024
Working-Age Population (PET) World Bank - WDI Annual 1991-2024
Labor Force (PEA) ILOSTAT Annual 1991-2024
Employment Rate Derived from levels Annual 1991-2024
Unemployment Rate Derived from levels Annual 1991-2024
Participation Rate Derived from levels Annual 1991-2024
Inactivity Rate Derived from levels Annual 1991-2024
Note: Author’s own construction.
Appendix J.2 Official National labor market indicators
Table J.6: National Labor and Demographic Indicators Used as Structural Baseline
Indicator Frequency Coverage
Total Population Annual 1993-2050
Working-Age Population (PET) Annual 1993-2050
Economically Active Population (PEA) Monthly 2001-2024
Employed Population Monthly 2001-2024
Unemployed Population Monthly 2001-2024
Inactive Population Monthly 2001-2024
Labor Participation Rate Monthly 2001-2024
Employment Rate Monthly 2001-2024
Unemployment Rate Monthly 2001-2024
Inactivity Rate Monthly 2001-2024
Note: Author’s own construction.
Appendix J.3 Official Departments Labor Market Indicators
Table J.7: Subnational Labor Indicators: Departments
Indicator Frequency Coverage
Employment Annual 33 departments (2005–2024)
Unemployment Annual 33 departments (2005–2024)
Economically Active Population (PEA) Annual 33 departments (2005–2024)
Working-Age Population (PET) Annual 33 departments (2005–2024)
Total Population Annual 33 departments (2005–2024)
Inactive Population Annual 33 departments (2005–2024)
Participation / Employment / Unemployment / Inactivity Rates Annual 33 departments (2005–2024)
Note: Author’s own construction.
Appendix J.4 Official Main Cities Labor Market Indicators
Table J.8: Subnational Labor Indicators: Main Cities
Indicator Frequency Coverage
Employment Monthly 32 cities + Cundinamarca (2021–2024)
Unemployment Monthly 32 cities + Cundinamarca (2021–2024)
Economically Active Population (PEA) Monthly 32 cities + Cundinamarca (2021–2024)
Working-Age Population (PET) Monthly 32 cities + Cundinamarca (2021–2024)
Total Population Monthly 32 cities + Cundinamarca (2021–2024)
Inactive Population Monthly 32 cities + Cundinamarca (2021–2024)
Participation / Employment / Unemployment / Inactivity Rates Monthly 32 cities + Cundinamarca (2021–2024)
Note: Author’s own construction.
Appendix J.5 Monthly Macroeconomic Indicators from Central Bank (Banco de la República)
Table J.9: Official Monthly Economic Indicators Used as Predictors
Indicator Frequency Coverage
Consumer Price Index (IPC, Índice de Precios al Consumidor) Monthly National (1991-2025)
Producer Price Index (IPP, Índice de Precios al Productor) Monthly National (1991-2025)
Representative Market Exchange Rate (TRM, Tasa Representativa del Mercado) Daily / Monthly National (1991-2025)
Monthly Inflation (variation of IPC) Monthly National (1991-2025)
Note: Author’s own construction.
Appendix J.6 Official Informality Indicators
Table J.10: Labor Informality Indicators: National and Cities
Indicator Frequency Coverage
Informality Rate (Methodology I) Monthly National total (2007–2020)
Informality Rate (Methodology II) Monthly National total (2021–Present)
Informality Rate (Methodology I) Monthly 23 main cities (2007–2020)
Informality Rate (Methodology II) Monthly 23 main cities (2021–Present)
Note: Author’s own construction.